### v1

# Mutation-acyclic quivers are totally proper

Abstract.
Totally proper quivers, introduced by S. Fomin and the author [11], have many useful properties including powerful mutation invariants. We show that every mutation-acyclic quiver (i.e., a quiver that is mutation equivalent to an acyclic one) is totally proper. This yields new necessary conditions for a quiver to be mutation-acyclic.
Key words and phrases:
Quiver, mutation, cyclic ordering, Alexander polynomial.
1991 Mathematics Subject Classification:
Primary 13F60, Secondary 05E16, 15B36.
Partially supported by NSF grants DMS-1840234, DMS-2054231, DMS-2348501.
1. Introduction
Quivers are finite directed graphs without oriented cycles of length $1$ or $2$. Mutations are operations that transform a quiver, based on a choice of a vertex. These notions are foundational in the study of cluster algebras [15]. A mutation invariant is a characteristic of a quiver that is preserved under mutations. Mutation invariants are helpful for deciding whether two quivers are mutation equivalent or not, i.e., whether there is a sequence of mutations that transforms one quiver into the other. See [5, 9, 14, 20, 22] for examples of known mutation invariants.
A cyclically ordered quiver (COQ) is a pair $(Q,\sigma)$ where $Q$ is a quiver and $\sigma$ a cyclic ordering of its vertices. Cyclically ordered quivers were introduced in [11] to develop new powerful mutation invariants. A mutation in a COQ $(Q,\sigma)$ transforms $Q$ by the usual mutation rule, while simultaneously changing the cyclic ordering $\sigma$ in a prescribed way. We note that mutations of COQs are only allowed at the vertices that satisfy a certain properness condition. It is this condition that ultimately enables the introduction of new mutation invariants.
In this paper, we prove that for one important class of quivers, the properness requirement can be lifted, so that mutations at all vertices are allowed and all invariants developed in [11] become true mutation invariants. Specifically, we show that in any mutation-acyclic quiver $Q$ (i.e., a quiver that is mutation equivalent to an acyclic quiver), every vertex $v$ is proper, for a particular canonical cyclic ordering $\sigma$ on $Q$. After a mutation at $v$, we obtain a new COQ $(Q^{\prime},\sigma^{\prime})$, with canonical cyclic ordering $\sigma^{\prime}$, that again has the same property: all the vertices are proper, so we can mutate at any one of them, and the process continues.
This theorem yields new mutation invariants of mutation-acyclic quivers, as well as new tools for proving that various quivers are not mutation-acyclic. One such quiver is shown in Figure 1. We give a short proof that this quiver is not mutation-acyclic in Example 6.11.
A graph with ten vertices and 18 arrows, arranged into an equalateral triangular grid. There are four vertices on each side, with one interior vertex, thus there are nine equalateral triangular faces (each supported by three vertices) which cover the whole. Each face has the arrows oriented cyclically, directed counter-clockwise.
Figure 1. An (unlabeled) triangular grid quiver with $4$ vertices on each side. Readers may recognize it as the default quiver in B. Keller’s mutation applet [16].
As mentioned above, mutation at a vertex $v$ in a COQ $(Q,\sigma)$ is only allowed if $v$ satisfies a combinatorial condition of “properness.” Informally, $v$ is proper if every 2-arrow oriented path through $v$ travels “clockwise,” i.e., in the direction of the cyclic ordering $\sigma$. A COQ is proper if every vertex $v$ in it is proper (possibly after applying a sequence of transpositions called wiggles). Thus, we can mutate at any vertex in a proper COQ $(Q,\sigma)$ and get a new COQ $(Q^{\prime},\sigma^{\prime})$—but $(Q^{\prime},\sigma^{\prime})$ will not necessarily be proper. If it is the case that applying any sequence of mutations to $(Q,\sigma)$ results in a proper COQ, then we say that $(Q,\sigma)$ is totally proper.
We have shown in [11, Theorem 11.3, Remark 11.11] that a quiver $Q$ can be upgraded to a totally proper COQ $(Q,\sigma)$ in at most one way (up to wiggles). Moreover, the (essentially unique) candidate cyclic ordering $\sigma=\sigma_{Q}$ can be constructed efficiently. We say that a quiver $Q$ is totally proper if the COQ $(Q,\sigma_{Q})$ is totally proper.
As shown in [11, Corollary 11.2], totally proper quivers have a powerful mutation invariant, which we recall next. This invariant is constructed as follows.
Input: a totally proper quiver $Q$.
Step 1. Construct the canonical cyclic ordering $\sigma=\sigma_{Q}$.
Step 2. “Tear” $\sigma$ into a linear ordering <<\,<.
Step 3. Construct the skew-symmetric exchange matrix $B=B_{Q}$, with rows and columns ordered acording to $<$.
Step 4. Construct the unipotent companion $U$, the unique unipotent upper-triangular matrix such that $B=U^{T}-U$.
Output: the integral congruence class of $U$, i.e., the set $\{GUG^{T}\mid G\in{\operatorname{GL}}_{n}({\mathbb{Z}})\}$.
We have shown in [11] that wiggles, cyclic shifts of the linear ordering, and (proper) mutations of COQs preserve the integral congruence class of the unipotent companion. While the resulting invariant of proper mutations is very powerful, it is not easy to use in practice, since the problem of deciding whether two upper-triangular matrices are congruent over ${\operatorname{GL}}_{n}({\mathbb{Z}})$ seems to be rather difficult. In [11], we bypassed this difficulty as follows. It is well known (and easy to see) that the integral congruence class of a matrix $U\in{\operatorname{GL}}_{n}({\mathbb{Z}})$ uniquely determines the conjugacy class of its cosquare $U^{-T}U$. It follows that whenever two COQs are related by proper mutations, the cosquares of their respective unipotent companions must be conjugate in ${\operatorname{GL}}_{n}({\mathbb{Z}})$. This conjugacy condition can be verified efficiently, though the algorithm is quite nontrivial [3, 7].
The cosquare $U^{-T}U$ can be used to construct other invariants of proper mutations, such as the monic characteristic polynomial of $U^{-T}U$, which we call the Alexander polynomial of the COQ $Q$. (While much simpler to compute, the Alexander polynomial is less powerful than the conjugacy class of $U^{-T}U$: integer matrices may have the same characteristic polynomial while not being conjugate in ${\operatorname{GL}}_{n}({\mathbb{Z}})$.)
The main result of this paper is the following theorem that settles [11, Conjecture 12.8].
Theorem 1.1.
Any mutation-acyclic quiver is totally proper.
Remark 1.2.
It suffices to show that any acyclic quiver is totally proper. Furthermore, as noted in [11, Observation 10.7], any acyclic quiver has a proper cyclic ordering obtained by “closing up” any linear ordering compatible with the orientations of the quivers’ arrows. (All such cyclic orderings are related by wiggles.) Our proof of Theorem 1.1 shows that these cyclic orderings are totally proper.
In the course of proving Theorem 1.1, we establish a few results that may be of independent interest. In particular, Theorem 4.4 gives a simple combinatorial constraint on the orientations of arrows in mutation-acyclic quivers. Both Theorem 4.4 and the ensuing Lemma 5.5 assert that certain subquivers of a mutation-acyclic quiver must be acyclic. Should someone wish to compute with COQs and unipotent companions, Lemmas 4.1 and 4.2 generalize and remove conditions from [11, Lemma 7.2 and Remark 7.3]. This can speed up and simplify the computations.
An alternative proof of Theorem 1.1 was independently discovered by Hugh Thomas, using categorification [24].
Overview of the paper
Section 2 establishes notation and reviews some material that does not involve COQs. Much of this section is devoted to restating the results of A. Seven [21], who constructed a distinguished symmetric matrix $A$ associated with an arbitrary mutation-acyclic quiver. Section 3 is a condensed summary of the necessary results and notation from [11]. It contains new examples, and Proposition 3.21 gives an alternative description of the Markov invariant. In Section 4 we establish a handful of technical lemmas. Lemmas 4.1 and 4.2 make it easier to compute the unipotent companion of a COQ after a proper mutation. Theorem 4.4 generalizes [11, Corollary 10.12] using the results of A. Seven [21]. Section 5 contains the proof of Theorem 1.1. We begin by defining $U=(A-B)/2$, where $A$ comes from Seven’s construction and $B$ is the usual exchange matrix. The fact that the integral congruence class of $U$ is a mutation invariant follows quickly from Proposition 2.30. The bulk of the effort is then spent showing that $U$ is indeed a unipotent companion. In Section 6, we give some applications and corollaries of Theorem 1.1. This includes a description of invariants for some small quivers, inequalities satisfied by the coefficients of the Alexander polynomial for mutation-acyclic quivers (Corollary 6.6), and several examples of quivers that are shown to be not mutation-acyclic.
Acknowledgements
I would like to thank my advisor, Sergey Fomin, for his guidance, comments, and suggestions; Grayson Moore, for reading and commenting on an early draft; and Danielle Ensign for her software assistance. I am also grateful to Roger Casals, Tucker Ervin, and Hugh Thomas for stimulating discussions.
I have used Magma and Sage for various computations. This research was supported in part through computational resources and services provided by Advanced Research Computing at the University of Michigan, Ann Arbor.
2. Quivers, mutation, and quasi-Cartan companions
We begin by reviewing notation, definitions and results from the literature which do not involve cyclically ordered quivers. This includes a summary of much from [21].
Definition 2.1.
A quiver is a finite directed graph with parallel edges allowed, but no directed $1$ or $2$-cycles. Directed edges in a quiver are called arrows. Each vertex is marked as mutable or frozen. Unless otherwise indicated, all vertices are mutable.
We use the notation $u\rightarrow v$ to assert that there is at least one arrow from $u$ to $v$. We use the notation $u\stackrel{{\scriptstyle x}}{{\rightarrow}}v$ to denote that there are $x$ arrows from vertex $u$ to vertex $v$. Let ${\operatorname{In}}(v)=\{u|u\rightarrow v\}$ and ${\operatorname{Out}}(v)=\{u|u\leftarrow v\}$ respectively denote the inset and outset of $v$ in a quiver.
Remark 2.2.
By default, our quivers have labeled vertices. Thus, we distinguish between isomorphic quivers on the same set of vertices that differ from each other by a permutation of this set.
Definition 2.3.
To mutate a quiver $Q$ at a mutable vertex $v$, do the following:
  1. (1)
    for each oriented path $u\rightarrow v\rightarrow w$, draw a new arrow $u\rightarrow w$ (thus, if there are $x$ arrows from $u\rightarrow v$ and $y$ arrows $v\rightarrow w$, we add $xy$ arrows $u\rightarrow w$);
  2. (2)
    reverse all arrows adjacent to $v$;
  3. (3)
    remove oriented $2$-cycles (one cycle at a time) until we again have a quiver.
We denote the resulting quiver by $\mu_{v}(Q)$. Mutation is an involution, $\mu_{v}(\mu_{v}(Q))=Q$. Mutation does not change which vertices are mutable or frozen.
Definition 2.4.
Two quivers are mutation equivalent if they are related by a sequence of mutations at mutable vertices. The mutation class $[Q]$ is the set of quivers mutation equivalent to $Q$.
Definition 2.5.
A quiver is acyclic if it does not contain any directed cycles. A quiver is mutation-acyclic if it is mutation equivalent to an acyclic quiver.
Example 2.6.
Only one of the two quivers in the Figure 2 is mutation-acyclic. Can you guess which is which?
Two cyclically oriented 3-vertex quivers, both with vertices labeled v1, v2 and v3. The cyclically oriented quiver on the left has multiplicities (weights) 5, 12, and 57 (between vertices 1 and 2, 2 and 3, and 3 and 1 respectively). The cyclically oriented quiver on the right has multiplicities 5, 12, and 58.
Figure 2. Two quivers on $3$-vertices. The quiver on the left is not mutation-acyclic, while the quiver on the right is mutation-acyclic (by mutating at $v_{2}$ and then $v_{1}$).
Definition 2.7.
For a quiver $Q$, the $B$-matrix $B_{Q}=(b_{uv})$ (also known as the exchange matrix) is the skew-symmetric adjacency matrix of $Q$. By convention we have $b_{uv}>0$ when there are $b_{uv}$ arrows $u\rightarrow v$, and $b_{uv}<0$ when there are $-b_{uv}$ arrows $u\leftarrow v$.
Definition 2.8.
A (full) subquiver is an induced subgraph of a quiver. We call the remaining vertices of a subquiver its support. The mutable part of a quiver $Q$ is the subquiver supported by its mutable vertices.
Definition 2.9.
The principal extension (also known as framing) of a quiver $Q$ is a quiver $\widehat{Q}$ formed by adding a new frozen vertex $v^{\prime}$ and a single arrow $v^{\prime}\stackrel{{\scriptstyle}}{{\rightarrow}}v$ for each vertex $v$. Thus $Q$ is the mutable part of $\widehat{Q}$ and $\widehat{Q}$ has twice as many vertices as $Q$.
The principle framing of an acyclic four vertex quiver, with vertices v1, v2, v3, and v4. There is one arrow from v1 to v2, one arrow from v1 to v3, three arrows from v1 to v4, two arrows from v2 to v4, and one arrow from v3 to v4. The four frozen vertices are denoted v1’, v2’, v3’, v4’, and each come with a single arrow from v i’ to v i.
Figure 3. The black circle vertices $v_{i}$ support an acyclic quiver $Q$. With the additional square (frozen) vertices $v_{i}^{\prime}$, we get the principal framing $\widehat{Q}$.
We will use the following notation for mutation classes and associated matrices and data. (Cf. also Definitions 2.24 and 5.1.)
Definition 2.10.
Let $\mathbb{T}_{n}$ be an $n$-regular tree whose edges are labeled by the integers $1,\ldots,n$, so that each edge label appears next to each vertex. We write $t\stackrel{{\scriptstyle i}}{{\text{---}}}t^{\prime}$ to indicate that an edge labeled $i$ joins the vertices $t$ and $t^{\prime}$ in $\mathbb{T}_{n}$. The tree $\mathbb{T}_{n}$ always comes with a distinguished vertex $t_{0}$.
Fix a quiver $Q_{0}$ with $n$ mutable linearly ordered vertices $v_{1}<\cdots<v_{n}$ (and no frozen vertices). We index the mutation class $[\widehat{Q}_{0}]$ by simultaneously assigning a quiver $\tilde{Q}_{t}\in[\widehat{Q}_{0}]$ to each vertex $t$ in $\mathbb{T}_{n}$, so that $\tilde{Q}_{t_{0}}=\widehat{Q}_{0}$ and $\tilde{Q}_{t}=\mu_{v_{i}}(\tilde{Q}_{t^{\prime}})$ whenever $t\stackrel{{\scriptstyle i}}{{\text{---}}}t^{\prime}$. We call $Q_{0}$ the initial quiver. For each vertex $t\in\mathbb{T}_{n}$, let $Q_{t}$ be the mutable part of $\tilde{Q}_{t}$. Let $B_{t}=(b_{v_{i}v_{j};t})$ be the $n\times n$ exchange matrix of $Q_{t}$. Similarly, let $C_{t}=(c_{v_{i}^{\prime}v_{j};t})$ be $n\times n$ bipartite adjacency matrix between frozen vertices and mutable vertices. Thus
$b_{v_{i}v_{j};t}=\#\{\text{arrows }v_{i}\rightarrow v_{j}\text{ in }Q_{t}\}-\#\{\text{arrows }v_{i}\leftarrow v_{j}\text{ in }Q_{t}\},$
$c_{v_{i}^{\prime}v_{j};t}=\#\{\text{arrows }v_{i}^{\prime}\rightarrow v_{j}\text{ in }\tilde{Q}_{t}\}-\#\{\text{arrows }v_{i}^{\prime}\leftarrow v_{j}\text{ in }\tilde{Q}_{t}\}.$
The matrix $C_{t}$ is called the $C$-matrix, and its columns $\mathbb{c}_{v_{i};t}$ are called $\mathbb{c}$-vectors [18]. Note that $B_{t}$ is skew-symmetric, while $C_{t}$ need not have any symmetries.
Whenever we use a quiver or matrix indexed by $t$ we will specify an initial quiver, or relevant conditions on that choice. We will often take an acyclic initial quiver $Q_{0}$.
Example 2.11.
Let $Q_{0}$ be the acyclic quiver with vertices $v_{1},v_{2},v_{3},v_{4}$, and arrows $v_{1}\stackrel{{\scriptstyle}}{{\rightarrow}}v_{2}\stackrel{{\scriptstyle 2}}{{\rightarrow}}v_{4}$, $v_{1}\stackrel{{\scriptstyle}}{{\rightarrow}}v_{3}\stackrel{{\scriptstyle}}{{\rightarrow}}v_{4}$, and $v_{1}\stackrel{{\scriptstyle 3}}{{\rightarrow}}v_{4}$ (see Figure 3 for $\widehat{Q}_{0}$). With initial quiver $Q_{0}$ (and linear order $v_{1}<v_{2}<v_{3}<v_{4}$), we have
$B_{t_{0}}=\begin{pmatrix}0&1&1&3\\ -1&0&0&2\\ -1&0&0&1\\ -3&-2&-1&0\end{pmatrix},\quad\quad C_{t_{0}}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\end{pmatrix}.$
Indeed, by construction $C_{t_{0}}$ is the identity matrix regardless of the initial quiver.
If $t\stackrel{{\scriptstyle 2}}{{\text{---}}}t_{0}$ then
$B_{t}=\begin{pmatrix}0&-1&1&5\\ 1&0&0&-2\\ -1&0&0&1\\ -5&2&-1&0\end{pmatrix},\quad\quad C_{t}=\begin{pmatrix}1&0&0&0\\ 0&-1&0&2\\ 0&0&1&0\\ 0&0&0&1\end{pmatrix}.$
Remark 2.12.
The map $t\mapsto(C_{t},B_{t})$ is called a (principal, tropical) $Y$-seed pattern in [21], as it is essentially equivalent to an assignment of tropical coefficient variables (usually denoted $\mathbb{y}$) and $B$-matrix to each vertex $t\in\mathbb{T}_{n}$, cf. [14, Proposition 3.6.5].
Definition 2.13.
A (mutable) vertex $v$ in a quiver $Q_{t}$ (or $\tilde{Q}_{t}$) is green (resp., red) if every entry of $\mathbb{c}_{v;t}$ is nonnegative (resp., nonpositive).
Example 2.14.
Every vertex in $Q_{t_{0}}$ is green. Mutation at vertex $v$ toggles the color of $v$. (Other vertices may change color as well!)
Theorem 2.15 ([6]).
Every vertex in $Q_{t}$ is either green or red (never both), regardless of the choice of initial quiver.
We next re-interpret Theorem 2.15 using the following definition.
Definition 2.16.
[…]
  • remove all frozen vertices and the arrows incident to them;
  • add a single new frozen vertex $\mathring{v}$;
  • for each green mutable vertex $v_{j}$, add $\sum_{i}c_{v_{i}^{\prime}v_{j};t}$ arrows $\mathring{v}\rightarrow v_{j}$;
  • for each red mutable vertex $v_{j}$, add $-\sum_{i}c_{v_{i}^{\prime}v_{j};t}$ arrows $\mathring{v}\leftarrow v_{j}$.
[…]
Corollary 2.17.
[…]
Remark 2.18.
More generally, $C$-matrices can be used to compute the number of arrows between frozen and mutable vertices for “triangular extensions” other than the principal framing, such as those that involve adding a frozen vertex $\mathring{v}$ and any number of arrows $\mathring{v}\rightarrow v$, cf. [4, Theorem 3.2].
Definition 2.19.
For a quiver $Q$ we define the underlying unoriented simple graph $K_{Q}$. The graph $K_{Q}$ has the same vertex set as $Q$, and an edge $u-v$ whenever there are any arrows $u\rightarrow v$ or $v\rightarrow u$ in $Q$. There are no parallel edges.
By a cycle in $K_{Q}$, we mean a sequence of vertices in $K_{Q}$ where each consecutive pair is joined by an edge $\mathcal{O}=(w_{0}-w_{1}-\cdots-w_{\ell}=w_{0})$, considered up to cyclic shifts. In particular, cycles in $K_{Q}$ have a direction of traversal. We will distinguish between two cycles with the same vertices traversed in the opposite order. Cycles thus correspond to an element of the first homology group $H_{1}(K_{Q},{\mathbb{Z}})$.
Definition 2.20 (cf. [11, Proposition 2.14]).
A cycle $\mathcal{O}=(w_{0}-w_{1}-\cdots-w_{\ell}=w_{0})$ in $K_{Q}$ is chordless if there are no edges between $w_{i}$ and $w_{j}$ for $i\neq j\pm 1$. If the arrows between $w_{i}$ and $w_{i+1}$ (in $Q$) are directed with the indexing of the cycle, $w_{i}\rightarrow w_{i+1}$, we say the cycle is forward-oriented. If instead the arrows are directed against the indexing, $w_{i}\leftarrow w_{i+1}$, then we say the cycle is backward-oriented. A cycle is oriented if it is either forward-oriented or backward-oriented.
Remark 2.21.
What we call a “chordless cycle in $K_{Q}$” is called a “cycle” in [21]. Thus, if a chordless cycle in $K_{Q}$ is oriented, then it is called an “oriented cycle.” Similarly, cycles which are not oriented are called ”nonoriented.”
In [21], cycles do not have an order of traversal, and so the forward/backward distinction does not arise. The order of traversal of cycles plays an important role in [11] (cf. Definition 3.8).
Two quivers on the vertices v1, v2, ..., v6. The quiver on the left is called Q, while the quiver on the right is called Q’. The quiver Q is an oriented cycles, with 3 arrows from v1 to v2 and one arrow from v2 to v3, v3 to v4, v4 to v5, v5 to v6, and from v6 to v1. There are no other arrows in Q. The quiver Q’ instead has 3 arrows from v1 to v2, one arrow from v3 to v2, v3 to v6, v3 to v4, v4 to v5, v5 to v6, and v6 to v1. Thus Q’ has an undirected cycle v1, v2, v3, v4, v5, v6, v1, but it has a chord from v3 to v6.
Figure 4. Two quivers $Q$ and $Q^{\prime}$. The cycle $(v_{1}-v_{2}-v_{3}-v_{4}-v_{5}-v_{6}-v_{1})$ in $K_{Q}$ is chordless and forward-oriented, while in $K_{Q^{\prime}}$ it is not chordless. The chordless cycles $(v_{1}-v_{2}-v_{3}-v_{6}-v_{1})$ and $(v_{3}-v_{4}-v_{5}-v_{6}-v_{3})$ in $K_{Q^{\prime}}$ are not oriented.
Definition 2.22 ([2]).
Fix a quiver $Q$ with $B$-matrix $B_{Q}=(b_{vw})$. A quasi-Cartan companion of $Q$ is a symmetric matrix $A=(a_{vw})$ such that $a_{vv}=2$ and $a_{vw}=\pm b_{vw}$ otherwise.
A quasi-Cartan companion of a given matrix $B_{Q}$ is determined by a choice of $\binom{n}{2}$ signs. The way these signs are chosen along chordless cycles will play a special role:
Definition 2.23 ([21]).
A quasi-Cartan companion $A=(a_{vw})$ of a quiver $Q$ is admissible if for every chordless cycle $\mathcal{O}=(w_{0}-\cdots-w_{\ell}=w_{0})$ in $K_{Q}$ the number #$\#\{i|a_{w_{i}w_{i+1}}>0\text{ and }i<\ell\}$ is odd when $\mathcal{O}$ is oriented, and is even otherwise.
Definition 2.24 (Cf. Definition 2.10).
Fix an acyclic initial quiver $Q_{0}$, with exchange matrix $B_{Q_{0}}=(b_{vw})$. We define the initial quasi-Cartan companion $A_{t_{0}}=(a_{vw})$ by $a_{vv}=2$ and $a_{vw}=-|b_{vw}|$ for $v\neq w$. For each $t\in\mathbb{T}_{n}$ we define the matrix $A_{t}=C_{t}^{T}A_{t_{0}}C_{t}$.
The following theorems of A. Seven will play an important role in our proof of Theorem 1.1.
Theorem 2.25 ([21, Theorem 1.3]).
Suppose the initial quiver $Q_{0}$ is acyclic. Then for every $t\in\mathbb{T}_{n}$, the matrix $A_{t}=C_{t}^{T}A_{t_{0}}C_{t}$ (see Definition 2.24) is an admissible quasi-Cartan companion of $Q_{t}$.
We will also need the following result.
Theorem 2.26 ([21, Theorems 1.3 and 1.4]).
Let $Q_{0}$ be acyclic, and fix $t\in\mathbb{T}_{n}$. Let $A_{t}=(a_{uv;t})$.
  • For an arrow $u\rightarrow v$ in $Q_{t}$, we have $a_{uv;t}>0$ if and only if $u$ is red and $v$ is green in $Q_{t}$.
  • Every oriented path of mutable vertices $w_{1}\rightarrow\cdots\rightarrow w_{m}$ in $Q_{t}$ has at most one positive entry in $A_{t}$ (thus $a_{w_{i}w_{i+1};t}>0$ for at most one $i<m$).
An illustration of Theorem 2.26 can be found in Example 2.32.
Remark 2.27.
To see how [21, Theorem 1.3] implies the first statement of Theorem 2.26, assume that $b_{ji}<0$ and check the four combinations of $\operatorname{sgn}(\mathbb{c}_{i})$ and $\operatorname{sgn}(\mathbb{c}_{j})$.
Remark 2.28.
For quivers mutation equivalent to acyclic quivers with sufficiently many arrows, a number of properties of the matrices $B_{t}$ and $C_{t}$ matrices have been recently established by T. Ervin, see [8, Proposition 5.2, Lemma 5.4]. These results may potentially be used to obtain enhancements of Theorem 2.26.
Definition 2.29 ([14, Theorem 2.8.3]).
We define a pair of square matrices $M_{Q}(v,\pm 1)$, associated to mutating a quiver $Q$ at a given vertex $v$. Let ${B_{Q}=(b_{pq})}$ be the exchange matrix of $Q$ and choose $\varepsilon=\pm 1$. Let $J$ be the diagonal matrix with $J_{vv}=-1$ and $J_{qq}=1$ for $w\neq v$. Let $E=(e_{pq})$ be the matrix with $v$th column $e_{qv}=\max(0,-\varepsilon b_{qv})$ and 00 otherwise. Then $M_{Q}(v,\varepsilon)=J+E$.
The matrix $M_{Q}$ can be used to mutate the $B$ and $C$-matrices (and sometimes the quasi-Cartan companion) associated with $Q$. Recall that red and green vertices are determined by the signs of the $C$-matrix (Definition 2.13).
Proposition 2.30 ( [14, Theorem 2.8.3], cf. [21, Lemma 3.1]).
Fix an initial quiver $Q_{0}$. Choose a quiver $\tilde{Q}_{t}\in[\widehat{Q}_{0}]$, and let $Q=Q_{t}$ (the mutable part of $\tilde{Q}_{t}$). Select a vertex $v_{i}$ of $Q$, and let $t\stackrel{{\scriptstyle i}}{{\text{---}}}t^{\prime}$. Then:
  • for either value of $\varepsilon=\pm 1$,
    (2.1) $B_{t^{\prime}}=M_{Q}(v_{i},\varepsilon)B_{t}M_{Q}(v_{i},\varepsilon)^{T};$
  • the $C$-matrices satisfy
    $C_{t^{\prime}}=\begin{cases}C_{t}M_{Q}(v_{i},1)^{T}&\text{if $v_{i}$ is green,}\\ C_{t}M_{Q}(v_{i},-1)^{T}&\text{if $v_{i}$ is red;}\end{cases}$
  • if the initial quiver $Q_{0}$ is acyclic then
    $A_{t^{\prime}}=\begin{cases}M_{Q}(v_{i},1)A_{t}M_{Q}(v_{i},1)^{T}&\text{if $v_{i}$ is green,}\\ M_{Q}(v_{i},-1)A_{t}M_{Q}(v_{i},-1)^{T}&\text{if $v_{i}$ is red.}\end{cases}$
The latter two formulas, involving $C_{t}$ and $A_{t}$, follow from their definitions and (2.1).
Example 2.31.
We demonstrate Proposition 2.30. Take our initial quiver $Q_{0}$ to be the acyclic quiver with vertices and arrows $v_{1}\stackrel{{\scriptstyle 2}}{{\rightarrow}}v_{3}\stackrel{{\scriptstyle}}{{\rightarrow}}v_{2}$ and $v_{1}\stackrel{{\scriptstyle}}{{\rightarrow}}v_{2}$. Let $t\in\mathbb{T}_{n}$ satisfy $t_{0}\stackrel{{\scriptstyle 3}}{{\text{---}}}t_{1}\stackrel{{\scriptstyle 1}}{{\text{---}}}t_{2}\stackrel{{\scriptstyle 2}}{{\text{---}}}t.$
Consider the matrices $B_{t},$ and $C_{t}$, and their associated $A_{t}$ shown below.
$B_{t}=\begin{pmatrix}0&3&-13\\ -3&0&5\\ 13&-5&0\end{pmatrix},C_{t}=\begin{pmatrix}8&-3&0\\ 3&-1&0\\ 3&-1&-1\end{pmatrix},$
$A_{t}=\begin{pmatrix}2&-3&13\\ -3&2&-5\\ 13&-5&2\end{pmatrix}=C_{t}^{T}\begin{pmatrix}2&-1&-2\\ -1&2&-1\\ -2&-1&2\end{pmatrix}C_{t}.$
By inspecting the signs of (the columns of) $C_{t}$, we see that $v_{1}$ is green while $v_{2}$ and $v_{3}$ are red in $Q_{t}$.
Let $t^{\prime}\stackrel{{\scriptstyle 1}}{{\text{---}}}t$ (so $Q_{t^{\prime}}=\mu_{v_{1}}(Q_{t})$). We compute the mutation at $v_{1}$ and find
$B_{t^{\prime}}=\begin{pmatrix}0&-3&13\\ 3&0&-34\\ -13&34&0\end{pmatrix}.$
The B-matrix $B_{t^{\prime}}$ for $\mu_{v_{1}}(Q_{t})$ can be computed from $B_{t}$ by performing either of two different congruences (corresponding to $M_{Q_{t}}(v_{1},1),$ and $M_{Q_{t}}(v_{1},-1)$ respectively):
$\begin{pmatrix}-1&0&0\\ 3&1&0\\ 0&0&1\end{pmatrix}B_{t}\begin{pmatrix}-1&0&0\\ 3&1&0\\ 0&0&1\end{pmatrix}^{T}=\begin{pmatrix}-1&0&0\\ 0&1&0\\ 13&0&1\end{pmatrix}B_{t}\begin{pmatrix}-1&0&0\\ 0&1&0\\ 13&0&1\end{pmatrix}^{T}=B_{t^{\prime}}.$
However the same operations applied to $A_{t}$ do not both result in quasi-Cartan companions. In this case, because $v_{1}$ is green, congruence with $M_{Q_{t}}(v_{1},1)$ gives the quasi-Cartan companion $A_{t^{\prime}}$:
$\begin{pmatrix}-1&0&0\\ 3&1&0\\ 0&0&1\end{pmatrix}\begin{pmatrix}2&-3&13\\ -3&2&-5\\ 13&-5&2\end{pmatrix}\begin{pmatrix}-1&0&0\\ 3&1&0\\ 0&0&1\end{pmatrix}^{T}=\begin{pmatrix}2&-3&-13\\ -3&2&34\\ -13&34&2\end{pmatrix}.$
If instead we multiply with the matrix $M_{Q_{t}}(v_{1},-1)$:
$\begin{pmatrix}-1&0&0\\ 0&1&0\\ 13&0&1\end{pmatrix}\begin{pmatrix}2&-3&13\\ -3&2&-5\\ 13&-5&2\end{pmatrix}\begin{pmatrix}-1&0&0\\ 0&1&0\\ 13&0&1\end{pmatrix}^{T}=\begin{pmatrix}2&3&-39\\ 3&2&-44\\ -39&-44&678\end{pmatrix}.$
Note in particular that one of the diagonal entries of the last matrix is not $2$.
If we instead consider $t^{\prime\prime}\stackrel{{\scriptstyle 3}}{{\text{---}}}t$, then the associated $B$ and $C$ matrices are:
$B_{t^{\prime\prime}}=\begin{pmatrix}0&-62&13\\ 62&0&-5\\ -13&5&0\end{pmatrix},C_{t^{\prime\prime}}=\begin{pmatrix}8&-3&0\\ 3&-1&0\\ 3&-6&1\end{pmatrix}.$
In this case, only congruence with $M_{Q_{t}}(v_{3},-1)$ gives a quasi-Cartan companion:
$\begin{pmatrix}1&0&0\\ 0&1&5\\ 0&0&-1\end{pmatrix}A_{t}\begin{pmatrix}1&0&0\\ 0&1&5\\ 0&0&-1\end{pmatrix}^{T}=\begin{pmatrix}2&62&-13\\ 62&2&-5\\ -13&-5&2\end{pmatrix}.$
Example 2.32.
Taking $A_{t},B_{t}$ and $Q_{t}$ as in Example 2.31, we see the only positive off-diagonal entry of $A_{t}$ is $a_{v_{1}v_{3};t}=13$. As predicted by Theorem 2.26 we see $v_{3}\rightarrow v_{1}$, $v_{3}$ is red, and $v_{1}$ is green in $Q_{t}$.
3. Cyclically ordered quivers
We recall many definitions, notations, and results from [11]. We include some new examples, but do not repeat proofs.
Definition 3.1.
A cyclic ordering of a set $V$ is a linear ordering considered up to cyclic shifts—that is, taking the minimal element $v$ and forming a new linear ordering where $v>u$ for all $u\neq v$ (the order is otherwise unchanged). By repeated cyclic shifts, any element can be made maximal or minimal. The cyclic ordering associated to a linear ordering $v_{1}<\cdots<v_{n}$ will be denoted $(v_{1},\ldots,v_{n})$.
Definition 3.2.
A cyclically ordered quiver $(Q,\sigma)$ (abbreviated COQ) is a quiver $Q$ equipped with a cyclic ordering of its (mutable) vertices $\sigma$. Likewise, a linearly ordered quiver is a quiver equipped with a linear ordering of its (mutable) vertices. We may omit the ordering when it is clear from context or not needed, and simply denote a COQ by $Q$. There are $n$ linearly ordered quivers associated to each COQ, given by “tearing” the cyclic order between different consecutive vertices in the cyclic ordering.
We say an oriented path $u\rightarrow v\rightarrow w$ makes a right turn at $v$ if $u<v<w$ in some linear order associated to $\sigma$.
Six distinct cyclically ordered quivers with the same underlying quiver, which is described in the caption. Each is drawn with the vertices placed on a dashed circle, with that order written below the cyclically ordered quiver. From left to right the orders are (v1, v2, v3, v4), (v1, v3, v2, v4), (v1, v2, v4, v3), (v1, v3, v4, v2), (v1, v4, v3, v2), (v1, v4, v2, v3).
Figure 5. The $4!/4=6$ distinct COQs on the $4$-vertex acyclic quiver with vertices and arrows $v_{1}\rightarrow v_{2}\stackrel{{\scriptstyle 2}}{{\rightarrow}}v_{4}$, $v_{1}\rightarrow v_{3}\rightarrow v_{4}$, and $v_{1}\stackrel{{\scriptstyle 3}}{{\rightarrow}}v_{4}$.
Definition 3.3.
The unipotent companion of a linearly ordered quiver $Q$ is the unipotent upper triangular matrix $U$ such that $U^{T}-U=B_{Q}$, where the rows and columns of $B_{Q}$ are written in the linear ordering. The matrix $U$ can be constructed by negating $B_{Q}$, then setting the lower triangular part to 00 and the diagonal entries to $1$. For a COQ $Q$, the unipotent companion of any associated linearly ordered quiver is a unipotent companion of $Q$.
Example 3.4.
Consider the leftmost COQ $Q$ depicted in Figure 5. This COQ has $4$ associated linearly ordered quivers. We list their respective unipotent companions and linear orderings:
$\begin{pmatrix}1&-1&-1&-3\\ 0&1&0&-2\\ 0&0&1&-1\\ 0&0&0&1\end{pmatrix}\quad\begin{pmatrix}1&0&-2&1\\ 0&1&-1&1\\ 0&0&1&3\\ 0&0&0&1\end{pmatrix}\quad\begin{pmatrix}1&-1&1&0\\ 0&1&3&2\\ 0&0&1&-1\\ 0&0&0&1\end{pmatrix}\quad\begin{pmatrix}1&3&2&1\\ 0&1&-1&-1\\ 0&0&1&0\\ 0&0&0&1\end{pmatrix}$
$\hskip 4.26773ptv_{1}<v_{2}<v_{3}<v_{4}\quad\quad v_{2}<v_{3}<v_{4}<v_{1}\hskip 17.07182ptv_{3}<v_{4}<v_{1}<v_{2}\quad\quad v_{4}<v_{1}<v_{2}<v_{3}$
Proposition 3.5 ([11, Proposition 4.4]).
All unipotent companions of a COQ are integrally congruent.
Recall that the integral congruence class of a unipotent companion $U$ is the set of matrices $\{GUG^{T}|G\in{\operatorname{GL}}_{n}({\mathbb{Z}})\}$.
Definition 3.6.
A wiggle is a transformation of a COQ which fixes the quiver $Q$, and transposes two adjacent vertices in the cyclic ordering that are not adjacent in $Q$. We say that two COQs are wiggle equivalent if they have the same quiver, and their cyclic orderings are related by a sequence of wiggles. The wiggle equivalence class of a COQ is the set of all wiggle equivalent COQs, see Figure 6. If $U$ is a unipotent companion of some COQ in a wiggle equivalence class, we say that $U$ is a unipotent companion of the class.
[…]
Figure 6. The $4$ wiggle equivalence classes of the COQs in Figure 5.
Proposition 3.7 ([11, Proposition 4.5]).
All unipotent companions of wiggle equivalent COQs are integrally congruent.
Definition 3.8 (Cf. [11, Definition 2.10]).
Fix a COQ $(Q,\sigma)$. Let $\mathcal{O}=(w_{0}-w_{1}-\cdots-w_{k}=w_{0})$ be a cycle in $K_{Q}$ (Definition 2.20). Fix a linear order $<$ associated to $\sigma$. The winding number ${\operatorname{wind}}_{\sigma}(\mathcal{O})$ is the (signed) number of times we “wrap around” the linear order:
(3.1) ${\operatorname{wind}}_{\sigma}(\mathcal{O})=\#\{w_{i}|w_{i}\rightarrow w_{i+1},w_{i}>w_{i+1}\}-\#\{w_{i}|w_{i}\leftarrow w_{i+1},w_{i}<w_{i+1}\}.$
We omit the (straightforward) proofs that ${\operatorname{wind}}_{\sigma}(\mathcal{O})$ does not depend on the choice of linear order associated to $\sigma$, and that this definition agrees with [11, Definition 2.10].
Example 3.9.
Let $Q$ be the underlying quiver of the COQs shown in Figure 6, and consider the (chordless) cycle $\mathcal{O}=(v_{1}-v_{2}-v_{4}-v_{1})$ in $K_{Q}$. The winding number of this undirected cycle is 00 in the three COQs in the top row and $1$ in the three COQs in the bottom row.
Remark 3.10.
The winding number corresponds to the winding number of certain maps from the cycle (as a cell complex) to the circle. Specifically, map the vertices onto the circle according to the cyclic ordering (not the order of traversal), and then map each edge to an arc between the endpoints, starting from the tail of an associated arrow and traveling clockwise to the head. See Figure 7.
On the left, a COQ with quiver whose arrows are v1 to v2, v2 to v4, v1 to v3, v3 to v4, and whose cyclic ordering is (v1, v2, v3, v4). On the right, the same vertices drawn on a dashed circle in the same positions as the COQ, but the arrows have been drawn along the dashed circle. The arrow from v1 to v2 now lies on the circular arc clockwise from the location of v1 (135 degrees on the circle) to the location of v2 (45 degrees). The arrow from v2 to v3 lies on the circular arc clockwise from the location of v2 to the location of v4 (-135 degrees). The arrow from v3 to v4 lies on the circular arc clockwise from the location of v3 (-45 degrees) to the location of v4. The arrow from v1 to v3 lies on the circular arc clockwise from the location of v1 to the location of v3. In particular, none of the arrows go through the point at 180 degrees, while they cover each point from 135 degrees to -135 degrees (traveling clockwise) twice.
Figure 7. A COQ $Q$ and the image of the cycle $\mathcal{O}=(v_{1}-v_{2}-v_{4}-v_{3}-v_{1})$ in $K_{Q}$, with the arrows drawn on the circle. In this case the image is contractible, and so the winding number is 00.
Theorem 3.11 ([11, Theorem 2.12]).
Two COQs $(Q,\sigma)$ and $(Q,\sigma^{\prime})$ are wiggle equivalent if and only if ${\operatorname{wind}}_{\sigma}(\mathcal{O})={\operatorname{wind}}_{\sigma^{\prime}}(\mathcal{O})$ for every undirected cycle $\mathcal{O}$ in $K_{Q}$.
We now turn to mutations of COQs.
Definition 3.12.
A vertex $v$ in a COQ $(Q,\sigma)$ is proper if every oriented path ${u\rightarrow v\rightarrow w}$ makes a right turn at $v$ (Definition 3.2). The wiggle equivalence class is proper if for every vertex $v$ there is a COQ in the wiggle equivalence class in which $v$ is proper.
Example 3.13.
In the wiggle equivalence classes shown in Figure 6, vertices $v_{1},v_{4}$ are proper vertices in each class, as they have no oriented paths through them. Vertex $v_{2}$ is proper only in the three COQs in the top row, while vertex $v_{3}$ is proper only in the three COQs in the leftmost column(s). Thus the wiggle equivalence class in the top left, using the cyclic orderings $(v_{1},v_{2},v_{3},v_{4})$ and $(v_{1},v_{3},v_{2},v_{4})$, is the only class which is proper.
Definition 3.14.
To mutate a COQ $Q$ at a proper vertex $v$, mutate the quiver as usual and reposition $v$ in the cyclic ordering so that $v$ is still proper (that is, move $v$ clockwise in the cyclic ordering past all the elements of ${\operatorname{In}}(v)$ in the mutated quiver, without passing any vertices in ${\operatorname{Out}}(v)$; all choices give wiggle equivalent COQs). We denote the resulting wiggle equivalence class by $\mu_{v}(Q)$. The proper mutation class of a COQ is the set of all COQs which can be obtained from the original by a sequence of proper mutations and wiggles.
A sequence of three proper mutations between four COQs. Starting on the left, the quiver has vertices v1, v2, v3, v4 (with that cyclic ordering), and an arrow from v2 to v1, v3 to v1, v3 to v4, as well as two arrows from v2 to v4 and 3 arrows from v4 to v1. In particular, v1 is a sink vertex. Thus there are no oriented paths through v1, and so v1 is proper. The first mutation is at v1, which only reverses the arrows touching v1, and leaves the cyclic ordering unchanged. Notice that the only oriented path of length two through v2 starts at v1 and ends at v4. This is a right turn, so v2 is proper. We mutate at v2 to arrive at a new COQ with cyclic ordering (v1, v3, v4, v2) and an arrow from v1 to v3, v3 to v4, v2 to v1, as well as five arrows from v1 to v4 and two arrows from v4 to v2. As v2 had an oriented path through it, we have changed its position in the cyclic ordering. We next wish to mutate at v3. It is proper as the only oriented 2-path goes from v1 to v4, and v1, v3, v4 is a right turn. Proper mutation at v3 results in the COQ with cyclic ordering (v1, v4, v3, v2) and a single arrow from v2 to v1, v3 to v1, v4 to v3, as well as two arrows from v4 to v2 and six arrows from v1 to v4.
Figure 8. A sequence of proper mutations.
Proposition 3.15 ([11, Proposition 6.4]).
Proper mutation is well defined up to wiggle equivalence—that is, if two COQs are wiggle equivalent and a vertex $v$ is proper in each, then performing the proper mutation $\mu_{v}$ to each results in the same wiggle equivalence class of COQs.
Theorem 3.16 ([11, Theorem 7.1]).
The unipotent companions of COQs which are related by a proper mutation are integrally congruent.
Definition 3.17 ([11, Definition 11.1]).
A (wiggle equivalence class of) COQ(s) is totally proper if every COQ in its proper mutation class is proper. We say a cyclic ordering $\sigma$ is totally proper for a quiver $Q$ if $(Q,\sigma)$ is totally proper. If $Q$ has a totally proper cyclic ordering, we say $Q$ is totally proper.
Certain unipotent companions are transformed by matrix conjugation with the same matrices (see Definition 2.29) that transform the $B$-matrix. One case was established in [11], and follows. We generalize this result in the next section.
Lemma 3.18 ([11, Lemma 7.2 and Remark 7.3]).
Suppose that $v_{j}$ is a proper vertex in the COQ $Q$. Let $U$ be a unipotent companion of $Q$ with linear order $<$ such that $v_{i}<v_{j}$ for $v_{i}\in{\operatorname{In}}(v_{j})$ and $v_{i}<v_{k}$ for $v_{k}\in{\operatorname{Out}}(v_{j})$. Then the matrix $M_{Q}(v_{j},-1)UM_{Q}(v_{j},-1)^{T}$ is a unipotent companion of $\mu_{v_{j}}(Q)$.
We conclude this section by recalling two easily expressed invariants of COQs, along with some examples. Definition 3.19 and Proposition 3.21 do not play any role in the proof of Theorem 1.1, but do appear prominently in applications and corollaries.
Definition 3.19 ([11][Definitions 8.1, 8.4]).
Given a COQ $Q$ with unipotent companion $U$, the Alexander polynomial of $Q$ is given by $\Delta_{Q}(t)=\det(tU-U^{T})$. The Markov invariant $M_{Q}$ is defined by $M_{Q}=n+(\text{coefficient of }t^{n-1}\text{ in }\Delta_{Q}(t))$.
Remark 3.20.
As the exchange matrix $B$ and unipotent companion $U$ of a linearly ordered quiver $Q$ are related by $B=U^{T}-U$, it follows that
$\det(B)=\det(U^{T}-U)=(-1)^{n}\Delta_{Q}(1).$
Proposition 3.21.
Let $Q$ be an $n$-vertex acyclic quiver. Fix a linear order $<$ of the vertices of $Q$ so that $v_{i}<v_{j}$ whenever $v_{i}\rightarrow v_{j}$. Let $U$ be the unipotent companion of the linearly ordered quiver $(Q,<)$. For a cycle $\mathcal{O}=(w_{0}-\cdots-w_{\ell}=w_{0})$ in $K_{Q}$, let $wt(\mathcal{O})=\prod_{i}|b_{w_{i}w_{i+1}}|$ if there is exactly one location $i$ with $w_{i}\leftarrow w_{i+1}$ (thus $w_{j}\rightarrow w_{j+1}$ for all $j\neq i$), and $wt(\mathcal{O})=0$ otherwise. Then the Markov invariant
$M_{Q}=\sum_{v_{i}<v_{j}}b_{v_{i}v_{j}}^{2}+\sum_{\mathcal{O}}wt(\mathcal{O}).$
Proof.
The $t^{n-1}$ coefficient of $\det(tU-U^{T})$ is negative of the trace of $U^{T}U^{-1}$. Let $N=I-U$, note $N$ is positive and strictly upper triangular. Thus
$U^{-1}=I+N+N^{2}+\cdots+N^{n-1},$
and so ${\operatorname{tr}}(U^{T}\!U^{-1})=\sum_{k=0}^{n}{\operatorname{tr}}(U^{T}\!N^{k})$. We compute ${\operatorname{tr}}(U^{T}\!I)\!=n$, ${\operatorname{tr}}(U^{T}\!N)\!=-\sum_{v_{i}<v_{j}}\!b_{v_{i}v_{j}}^{2}$ and ${\operatorname{tr}}(U^{T}\!N^{k})=-\sum_{\mathcal{O}}wt(\mathcal{O}),$ where the sum is over the cycles $\mathcal{O}$ of length $k$. ∎
Example 3.22.
Let $Q$ be the leftmost COQ shown in Figure 5. Then
$\Delta_{Q}(t)=t^{4}+21t^{3}-43t^{2}+21t+1,$
and $M_{Q}=1^{2}+1^{2}+1^{2}+2^{2}+3^{2}+1\cdot 2\cdot 3+1\cdot 1\cdot 3=4+21=25.$
Remark 3.23.
Amanda Schwartz has recently developed a combinatorial description for all of the coefficients of the Alexander polynomial $\Delta_{Q}(t)$ in the case that $Q$ is a tree quiver (that is, $K_{Q}$ is a tree) [19].
Further examples appear in Section 6.
4. Technical Lemmas
In this section we establish lemmas and propositions which we will need in the proof of Theorem 1.1. We begin with generalizations of Lemma 3.18.
Lemma 4.1.
Suppose that $v_{j}$ is a proper vertex in the COQ $Q$. Let $U=(u_{v_{i}v_{k}})$ be a unipotent companion of $Q$ with linear order $<$ such that $v_{i}<v_{j}$ for $v_{i}\in{\operatorname{In}}(v_{j})$. Then the matrix $M_{Q}(v_{j},-1)UM_{Q}(v_{j},-1)^{T}$ is a unipotent companion of $\mu_{v_{j}}(Q)$.
Proof.
We follow the proof of [11, Lemma 7.2] closely. Throughout ${\operatorname{In}}(v_{j})$ and ${\operatorname{Out}}(v_{j})$ respectively refer to the inset and outset of $v_{j}$ in $Q$.
For the mutated quiver $Q^{\prime}=\mu_{v_{j}}(Q)$ let $<^{\prime}$ be a linear order such that
  • $v_{j}<^{\prime}v_{i}$ for all $v_{i}\in{\operatorname{In}}(v_{j})$,
  • if both $v_{k}\in{\operatorname{Out}}(v_{j})$ and $v_{k}<v_{j}$ then $v_{k}<^{\prime}v_{j}$, and
  • otherwise $<^{\prime}$ agrees with $<$ (so we have only changed the relative position of $v_{j}$).
Such an order exists because $v_{j}$ is proper in $Q$. Let $U^{\prime}=(u^{\prime}_{v_{i}v_{k}})$ be the unipotent companion of $Q^{\prime}$ with linear ordering $<^{\prime}$.
Let $B_{Q}=(b_{v_{i}v_{k}})$, set $n_{v_{i}v_{j}}=-\max(0,b_{v_{i}v_{j}})$ and also
$\delta_{v_{i}}=\begin{cases}-1&\text{if }v_{i}=v_{j};\\ 1&\text{else.}\end{cases}$
We wish to show that:
(4.1) $u^{\prime}_{v_{i}v_{k}}=\delta_{v_{i}}\delta_{v_{k}}u_{v_{i}v_{k}}-n_{v_{i}v_{j}}u_{v_{j}v_{k}}\delta_{v_{k}}-\delta_{v_{i}}u_{v_{i}v_{j}}n_{v_{k}v_{j}}+n_{v_{i}v_{j}}n_{v_{k}v_{j}}.$
Let $u^{\prime\prime}_{v_{i}v_{k}}$ denote the right hand side of (4.1).
By construction, we have
$u^{\prime}_{v_{i}v_{k}}=\begin{cases}1&\text{if }v_{i}=v_{k};\\ b_{v_{j}v_{k}}&\text{if }v_{i}=v_{j}<^{\prime}v_{k};\\ b_{v_{i}v_{j}}&\text{if }v_{i}<^{\prime}v_{j}=v_{k};\\ -b_{v_{i}v_{k}}-n_{v_{i}v_{j}}n_{v_{k}v_{j}}+\max(0,-b_{v_{i}v_{j}})\max(0,-b_{v_{j}v_{k}})&\text{if }v_{j}\neq v_{i}<^{\prime}v_{k}\neq v_{j};\\ 0&\text{if }v_{i}>^{\prime}v_{k}.\end{cases}$
Every case where both $v_{j}\leq^{\prime}v_{i},$ and $v_{j}\leq^{\prime}v_{k}$ appears in [11, Lemma 7.2], and the arguments therein apply. We consider the cases where $v_{j}$ appears after $v_{i}$ or $v_{k}$ in $<^{\prime}$. Note by construction of $<^{\prime}$ that if $w<^{\prime}v_{j}$ then $u_{v_{j}w}=0=n_{wv_{j}}$ and $b_{wv_{j}}=-u_{wv_{j}}.$
  • If $v_{i}=v_{k}<^{\prime}v_{j}$ then $u^{\prime}_{v_{i}v_{i}}=1-0-0+0=u^{\prime\prime}_{v_{i}v_{i}}.$
  • If $v_{i}<^{\prime}v_{j}=v_{k}$ then, $u^{\prime}_{v_{i}v_{j}}=b_{v_{i}v_{j}}=-u_{v_{i}v_{j}}-0-0+0=u^{\prime\prime}_{v_{i}v_{j}}$.
  • We split $v_{j}\neq v_{i}<^{\prime}v_{k}\neq v_{j}$ into $2$ cases, depending on the relative position of $v_{j}$.
    • If $v_{i}<^{\prime}v_{j}<^{\prime}v_{k}$ then $n_{v_{k}v_{j}}=-\max(0,b_{v_{k}v_{j}})$ and $u_{v_{i}v_{j}}=\max(0,-b_{v_{i}v_{j}})$. So
      $u^{\prime}_{v_{i}v_{k}}=-b_{v_{i}v_{k}}+\max(0,-b_{v_{i}v_{j}})\max(0,-b_{v_{j}v_{k}})=u_{v_{i}v_{k}}-0-u_{v_{i}v_{j}}n_{v_{k}v_{j}}+0=u^{\prime\prime}_{v_{i}v_{k}}.$
    • If $v_{i}<^{\prime}v_{k}<^{\prime}v_{j}$, then $u^{\prime}_{v_{i}v_{k}}=-b_{v_{i}v_{k}}=u_{v_{i}v_{k}}-0-0+0=u^{\prime\prime}_{v_{i}v_{k}}.$
  • If $v_{k}<^{\prime}v_{i}$ and $v_{k}<^{\prime}v_{j}$ then $u^{\prime}_{v_{i}v_{k}}=0-0-0+0=u^{\prime\prime}_{v_{i}v_{k}}.$
Let $\pi$ be the permutation matrix which transforms $<$ into $<^{\prime}$. Equation (4.1) can be rewritten as $U^{\prime}=\pi^{T}M_{Q}(v_{j},-1)UM_{Q}(v_{j},-1)^{T}\pi$ (recall the matrices $E$ and $J$ defined in Definition 2.29). The claim follows. ∎
We establish a corresponding version of Lemma 4.1 for $\varepsilon=1$.
Lemma 4.2.
Suppose that $v_{j}$ is a proper vertex in the COQ $Q$. Let $U$ be a unipotent companion of $Q$ with linear order $<$ such that $v_{j}<v_{k}$ for $v_{k}\in{\operatorname{Out}}(v_{j})$. Then the matrix $M_{Q}(v_{j},1)UM_{Q}(v_{j},1)^{T}$ is a unipotent companion of $\mu_{v_{j}}(Q)$.
Proof.
We reduce to Lemma 4.1 by introducing the opposite linearly ordered quiver $R^{op}$ of another linearly ordered $R$. The quiver of $R^{op}$ has the same vertices as $R$, but each arrow is reversed ($v\rightarrow w$ becomes $v\leftarrow w$) and the linear ordering is reversed ($v<w$ becomes $w<v$). It is easy to see that if $U_{R}$ is a unipotent companion of $R$, then $U_{R}^{T}$ is a unipotent companion of $R^{op}$ (note that $U_{R}^{T}$ is upper triangular in the reversed linear order). We also have $M_{R}(v_{j},1)=M_{R^{{op}}}(v_{j},-1)$.
Now we compute:
$M_{Q}(v_{j},1)UM_{Q}(v_{j},1)^{T}$ $=(M_{Q}(v_{j},1)U^{T}\!M_{Q}(v_{j},1)^{T})^{T}\!$
$=(M_{Q^{op}}(v_{j},-1)U^{T}\!M_{Q^{op}}(v_{j},-1)^{T})^{T}.$
Lemma 4.1 implies that $M_{Q^{op}}(v_{j},-1)U^{T}M_{Q^{op}}(v_{j},-1)^{T}$ is a unipotent companion of $Q^{op}$. Thus its transpose is a unipotent companion of $Q$, as desired. ∎
Example 4.3.
Let $Q$ be the leftmost COQ depicted in Figure 5. The unipotent companions of $Q$ appear in Example 3.4, let $U$ be the rightmost unipotent companion corresponding to the linear order $v_{4}<v_{1}<v_{2}<v_{3}$. Vertex $v_{2}$ is proper (and both $Q$ and $\mu_{v_{2}}(Q)$ appear in the center of Figure 8). We compute the unipotent companion $M_{Q}(v_{2},-1)UM_{Q}(v_{2},-1)^{T}$ of $\mu_{v_{2}}(Q)$:
$\begin{pmatrix}1&0&0&0\\ 0&1&1&0\\ 0&0&-1&0\\ 0&0&0&1\end{pmatrix}\begin{pmatrix}1&3&2&1\\ 0&1&-1&-1\\ 0&0&1&0\\ 0&0&0&1\end{pmatrix}\begin{pmatrix}1&0&0&0\\ 0&1&1&0\\ 0&0&-1&0\\ 0&0&0&1\end{pmatrix}^{T}$ $=\begin{pmatrix}1&5&-2&1\\ 0&1&0&-1\\ 0&-1&1&0\\ 0&0&0&1\end{pmatrix}.$
This matrix is upper triangular with the linear order $v_{4}<v_{2}<v_{1}<v_{3}$ (thus we have swapped the $2$nd and $3$rd rows and columns), in agreement with Lemma 4.1.
If we instead take $U^{\prime}$ to be the leftmost unipotent companion depicted in Example 3.4 (linear order $v_{1}<v_{2}<v_{3}<v_{4}$), then we may apply Lemma 4.2. We find another unipotent companion $M_{Q}(v,1)U^{\prime}M_{Q}(v,1)^{T}$:
$\begin{pmatrix}1&0&0&0\\ 0&-1&0&0\\ 0&0&1&0\\ 0&2&0&1\end{pmatrix}\begin{pmatrix}1&-1&-1&-3\\ 0&1&0&-2\\ 0&0&1&-1\\ 0&0&0&1\end{pmatrix}\begin{pmatrix}1&0&0&0\\ 0&-1&0&0\\ 0&0&1&0\\ 0&2&0&1\end{pmatrix}^{T}$ $=\begin{pmatrix}1&1&-1&-5\\ 0&1&0&0\\ 0&0&1&-1\\ 0&-2&0&1\end{pmatrix}.$
This matrix is upper triangular with order $v_{1}<v_{3}<v_{4}<v_{2}$.
We next establish that certain subquivers in a mutation-acyclic quiver are acyclic.
Theorem 4.4.
Suppose $Q$ is mutation-acyclic, and $v$ is a vertex in $Q$. Then the subquiver of $Q$ supported by ${\operatorname{Out}}(v)$ (resp. ${\operatorname{In}}(v)$) is acyclic.
This theorem generalizes an example provided in [21, Figure 1].
Proof.
As $Q$ is mutation-acyclic, by Theorem 2.25, $Q$ has an admissible quasi-Cartan companion $A=(a_{pq})$ (Definition 2.23). Thus every subquiver of $Q$ also has an admissible quasi-Cartan companion. (By restricting the rows and columns of $A$.)
Thus we may reduce to the case that $v$ is a source (resp. sink), and $v$ is adjacent to all other vertices in $Q$. If ${\operatorname{Out}}(v)$ (resp. ${\operatorname{In}}(v)$) contains an oriented cycle then it also contains an oriented chordless cycle (by a standard subdivision argument, any chord results in a smaller oriented cycle; cf. Lemma 4.6).
Assume for contradiction that the undirected simple graph $K_{Q}$ consists only of a chordless forward-oriented cycle $\mathcal{O}=(w_{0}-\cdots-w_{\ell}=w_{0})$ and the source vertex $v$ with edges $w_{i}-v$ (the case with $v$ a sink is identical). Then an odd number of the $a_{w_{i}w_{i+1}}$ are positive. In particular, at least one $a_{w_{i}w_{i+1}}$ is positive.
Without loss of generality, say $a_{w_{0}w_{1}}>0$. Note that $A$ remains an admissible quasi-Cartan companion if we simultaneously toggle the signs of the row and column corresponding to $w_{i}$. By repeatedly toggling signs, we may assume that $a_{w_{i}w_{i+1}}<0$ for all $i>0$ [why?]. However, each of the chordless cycles $(w_{i}-w_{i+1}-v-w_{i})$ in $K_{Q}$ are not oriented, and so any admissible companion must satisfy an even number of the inequalities ${a_{w_{i}w_{i+1}}>0},{a_{w_{i+1}v}>0},$ and ${a_{w_{i}v}>0}$. In particular, one of $a_{w_{0}v}$ or $a_{w_{1}v}$ is positive, without loss of generality say $a_{w_{1}v}>0$. As $a_{w_{1}w_{2}}<0$, we must have $a_{w_{2}v}>0$ for the chordless cycle $(w_{1}-w_{2}-v-w_{1})$ to have an even number of positive signs, and by induction we find that $a_{w_{i}v}>0$ for all $i$. But then the chordless cycle $(w_{0}-w_{1}-v-w_{0})$ has three positive entries in $A$, a contradiction. ∎
Remark 4.5.
Suppose $Q$ is mutation-acyclic, and $v$ is a vertex in $Q$. Partition the other vertices of $Q$ into three sets ${\operatorname{Out}}(v),{\operatorname{In}}(v)$ and the set of vertices $S$ which are not adjacent to $v$. Consider the subquivers supported by ${\operatorname{Out}}(v)\cup S$ (resp. ${\operatorname{In}}(v)\cup S$). Essentially the same argument as in Theorem 4.4 implies that every oriented cycle in these subquivers contains at most one vertex (and thus no arrows) in ${\operatorname{Out}}(v)$ (resp. ${\operatorname{In}}(v)$). Thus, for example, the quiver in Figure 9 is not mutation-acyclic.
A six vertex quiver, consisting of an oriented five-cycle v1, v2, v3, v4, v5, v1 and a vertex v6 with arrows from v1 to v6 and v4 to v6.
There is a single arrow between v2 and v3, v3 and v4, v4 and v5, v5 and v1, as well as three arrows from v1 to v2. Note that the number of arrows is irrelevant for the example; what matters is the orientation of the arrows and which vertices are in the inset or outset.
Figure 9. A quiver on $6$ vertices which does not have an admissible quasi-Cartan companion, and so cannot be mutation-acyclic. The subquiver supported by $v_{1},\ldots,v_{5}$ is an oriented cycle, contains two vertices in ${\operatorname{Out}}(v_{6})$, but no vertices in ${\operatorname{In}}(v_{6})$.
The following elementary lemma lets us convert a cycle in the underlying unoriented graph (Definition 2.19) into a chordless cycle, while keeping a given vertex on an oriented path.
Lemma 4.6.
Fix a vertex $v$ in a quiver $Q$. Let $\mathcal{O}$ be a cycle in $K_{Q}$, which contains the path $u\rightarrow v\rightarrow w$ for some vertices $u,w$. Then there exists a chordless cycle that is supported by a subset of the vertices of $\mathcal{O}$ and contains a path $u^{\prime}\rightarrow v\rightarrow w^{\prime}$ for some vertices $u^{\prime},w^{\prime}$, not necessarily distinct from $u,w$.
Proof.
We induct on the size of $\mathcal{O}$. If $\mathcal{O}$ is chordless (i.e., if $\mathcal{O}$ only contains $u,v,$ and $w$), then there is nothing to prove. Otherwise, $\mathcal{O}$ contains a chord. If $\mathcal{O}$ contains a chord which does not involve $v$, then we get a shorter chordless cycle by including the chord edge, and removing vertices and edges so that $u,v,$ and $w$ remain. Suppose instead that $\mathcal{O}$ contains a chord with the vertex $v$, without loss of generality say $v\rightarrow w^{\prime}$ for a vertex $w^{\prime}$ in $\mathcal{O}$ and distinct from $u,w$. We can replace the edge $v\rightarrow w$ by $v\rightarrow w^{\prime}$ and remove vertices so that $u,v$ and $w^{\prime}$ remain (and $w$ does not). The claim follows by induction. ∎
Finally, we give a new way to check that two COQs are wiggle equivalent.
Lemma 4.7.
Fix a quiver $Q$ and two linear orderings of its vertices <,<,< , and $<^{\prime}$. Let $U=(u_{vw})$ and $U^{\prime}=(u^{\prime}_{vw})$ be the unipotent companions of the linearly ordered quivers $(Q,<)$ and $(Q,<^{\prime})$ respectively. If $u_{vw}=u^{\prime}_{vw}$ for every pair of vertices $v,w$, then the COQs associated to $(Q,<)$ and $(Q,<^{\prime})$ are wiggle equivalent.
Remark 4.8.
Note that the condition $u_{vw}=u^{\prime}_{vw}$ for each pair of vertices is not the same as requiring that the matrices $U$ and $U^{\prime}$ are equal when they are written in their own respective linear orderings. For example, the vertex $v$ in the first subscript may refer to a different row than in the second.
Proof of Lemma 4.7.
By Theorem 3.11, it suffices to check that both COQs have the same winding numbers. From (3.1), it suffices to identify the arrows which “wrap around” the respective linear orderings (and, if they differ, their orientations). For any two vertices $v,w$ we have $v<w$ if and only if $u_{vw}\neq 0$. Thus the same arrows contribute regardless of if we use the linear ordering $<$ or $<^{\prime}$, and so the winding numbers agree. ∎
Example 4.9.
Let $Q$ be the acyclic quiver described in Figure 5. Consider the two unipotent companions from different linear orderings of $Q$ below.
$U=\begin{pmatrix}1&3&2&1\\ 0&1&-1&-1\\ 0&0&1&0\\ 0&0&0&1\end{pmatrix}\quad U^{\prime}=\begin{pmatrix}1&3&1&2\\ 0&1&-1&-1\\ 0&0&1&0\\ 0&0&0&1\end{pmatrix}$
$\quad\quad v_{4}<v_{1}<v_{2}<v_{3}\quad\quad\quad\quad v_{4}<v_{1}<v_{3}<v_{2}$
Each entry of $U$ matches $U^{\prime}$; for example, the $2$ in the first row and third column of $U$ corresponds to $u_{v_{4},v_{2}}$ and agrees with the entry in the first row and fourth column of $U^{\prime}$. Thus $U$ and $U^{\prime}$ are wiggle equivalent, indeed they are related by the wiggle $(v_{2}v_{3})$.
5. Proof of Theorem 1.1
We begin by establishing and recalling notation for this section. Using the notation from Definition 2.10, suppose our initial quiver $Q_{0}$ is acyclic with (mutable) vertices $v_{1},\ldots,v_{n}$. Fix a quiver $\tilde{Q}_{t}\in[\widehat{Q}_{0}]$, with mutable part $Q_{t}$. Thus $C_{t}$ denotes the $C$-matrix of $Q_{t}$, and $B_{t}$ denotes the $B$-matrix of $Q_{t}$. Finally, recall from Definition 2.24 the associated quasi-Cartan companions $A_{t_{0}}$ and $A_{t}=(a_{ij;t})=C_{t}^{T}A_{t_{0}}C_{t}$.
Definition 5.1.
We associate to each $t\in\mathbb{T}_{n}$ an additional matrix
$U_{t}=(u_{vw;t})=(A_{t}-B_{t})/2.$
Observe that $u_{vv;t}=1$ and $U_{t}^{T}-U_{t}=B_{t}$.
Example 5.2.
Let $Q_{0}$ be the acyclic quiver of type $E_{6}$ shown (as a COQ) in Figure 10, and let $t\in\mathbb{T}_{n}$ be the vertex such that
$t_{0}\stackrel{{\scriptstyle 3}}{{\text{---}}}t_{1}\stackrel{{\scriptstyle 2}}{{\text{---}}}t_{2}\stackrel{{\scriptstyle 5}}{{\text{---}}}t_{3}\stackrel{{\scriptstyle 4}}{{\text{---}}}t_{4}\stackrel{{\scriptstyle 2}}{{\text{---}}}t_{5}\stackrel{{\scriptstyle 1}}{{\text{---}}}t_{6}=t.$
In this case,
$B_{t}=\begin{pmatrix}0&0&0&1&0&-1\\ 0&0&-1&-1&1&1\\ 0&1&0&0&-1&-1\\ -1&1&0&0&0&0\\ 0&-1&1&0&0&0\\ 1&-1&1&0&0&0\end{pmatrix},\quad C_{t}=\begin{pmatrix}-1&0&0&0&0&1\\ 0&0&0&-1&0&1\\ 0&-1&0&0&0&1\\ 0&-1&1&0&0&0\\ 0&0&0&0&-1&0\\ 0&0&0&0&0&1\end{pmatrix},$
and therefore
$A_{t}=\begin{pmatrix}2&0&0&-1&0&-1\\ 0&2&-1&-1&-1&1\\ 0&-1&2&0&1&-1\\ -1&-1&0&2&0&0\\ 0&-1&1&0&2&0\\ -1&1&-1&0&0&2\end{pmatrix},\quad U_{t}=\begin{pmatrix}1&0&0&-1&0&0\\ 0&1&0&0&-1&0\\ 0&-1&1&0&1&0\\ 0&-1&0&1&0&0\\ 0&0&0&0&1&0\\ -1&1&-1&0&0&1\end{pmatrix}.$
All of the above is written using the linear order $v_{1}<v_{2}<\cdots<v_{6}$.
Two quivers on six vertices v1, v2, v3, v4, v5, v6. On the left, a tree quiver with an arrow from v1 to v2, v2 to v3, v3 to v4, v4 to v5, and v3 to v6. On the right, we have an arrow from v1 to v4, v4 to v2, v2 to v6, v6 to v1, v6 to v3, v3 to v2, v2 to v5, and v5 to v3.
$Q_{0}$        $Q_{t}$
Figure 10. The quiver $Q_{0}$ of type $E_{6}$, and a mutation equivalent quiver $Q_{t}$.
We will construct a family of linear orderings such that $U_{t}$ is upper triangular when written using one of them (see Lemma 5.8). To facilitate the construction, we first create several new directed graphs from $Q_{t}$.
Definition 5.3.
Fix a vertex $v_{j}$ of $Q_{t}$. Let $L_{v_{j}}(Q_{t})$ be the directed graph obtained by performing the following operations on $Q_{t}$:
  1. (1)
    if $v_{j}$ is red (resp., green), then for each oriented $2$ path $v_{i}\rightarrow v_{j}\rightarrow v_{k}$ with $v_{i},v_{k}$ both green (resp., red) in $Q_{t}$, add an arrow $v_{k}\stackrel{{\scriptstyle v_{j}}}{{\rightarrow}}v_{i}$ labeled with the vertex $v_{j}$;
  2. (2)
    reverse all arrows from a red vertex directed toward a green vertex. Equivalently, by Theorem 2.26, reverse all arrows $v\rightarrow w$ with $a_{vw;t}>0$.
Example 5.4.
Continuing with Example 5.2, we see from the columns of $C_{t}$ that $v_{3},v_{6}$ are green while $v_{1},v_{2},v_{4},$ and $v_{5}$ are red in $Q_{t}$. Therefore each $L_{v_{i}}(Q_{t})$ will reverse the arrows $v_{2}\rightarrow v_{6}$ and $v_{5}\rightarrow v_{3}$ (agreeing with the two positive off-diagonal entries of $A_{t}$), and we only add a new arrow to $L_{v_{i}}(Q_{t})$ if $i\in\{2,3,6\}$.
The directed graphs associated to vertices v1, v2, and v3. In each, vertices v3 and v6 are colored green, while v1, v2, v4, and v5 are colored red. Each has six vertices v1, v2, v3, v4, v5, v6, and all have edges from v1 to v4, v4 to v2, v2 to v5, v6 to v3, v6 to v2, v6 to v1, v3 to v2, and v3 to v5. The graph associated to v1 has no other edges. The graph associated to v2 has an additional edge from v6 to v3 which is labeled v2. The graph associated to v3 has an additional edge from v2 to v5 labeled v3.
The graphs associated to vertices v4, v5, and v6. The graphs have the same vertices and arrows as v1 in the previous figure, with the only additional edge appearing in the graph associated to v6. The new edge is from v1 to v2, and is labeled v6.
Figure 11. The six directed graphs $L_{v_{i}}(Q_{t})$ associated to $Q_{t}$ in Example 5.2. Vertices $v_{3},v_{6}$ are green, while vertices $v_{1},v_{2},v_{4},v_{5}$ are red.
Lemma 5.5.
Each directed graph $L_{v}(Q_{t})$ is acyclic. (In particular, there are no oriented $2$-cycles; $L_{v}(Q_{t})$ is an acyclic quiver with some marked arrows.) Furthermore, the subquiver of $Q_{t}$ supported by all red (resp., green) vertices is acyclic.
Proof.
By construction, there are no directed paths from a red vertex to a green vertex in $L_{v}(Q_{t})$. So any oriented cycle in $L_{v}(Q_{t})$ must have all vertices of the same color (red or green).
[…]
Without loss of generality, suppose $v$ is red. The case where $v$ is green is similar. Then we do not add any arrows labeled $v$ between red vertices, so we have shown there are no oriented cycles of red vertices. Further, any directed cycle involving green vertices must use at least one arrow labeled $v$ (those created in Step $1$ above). As all arrows labeled $v$ in $L_{v}(Q_{t})$ are oriented from ${\operatorname{Out}}(v)$ towards ${\operatorname{In}}(v)$, any oriented cycle must also involve at least one arrow which is not labeled $v$. We will now argue that any other cycles in $L_{v}(Q_{t})$ imply a problematic cycle $\mathcal{O}$ in the underlying unoriented simple graph $K_{Q_{t}}$ of $Q_{t}$.
Suppose there is an oriented cycle $w_{0}\stackrel{{\scriptstyle v}}{{\rightarrow}}w_{1}\rightarrow\cdots\rightarrow w_{k}=w_{0}$ in $L_{v}(Q_{t})$ that involves just one arrow $w_{0}\stackrel{{\scriptstyle v}}{{\rightarrow}}w_{1}$ labeled $v$. By construction of $L_{v}(Q_{t})$, there is a path ${w_{0}\leftarrow v\leftarrow w_{1}}$ in $Q_{t}$. Thus there is a cycle ${\mathcal{O}=(w_{0}-v-w_{1}-w_{2}-\cdots-w_{k})}$ in $K_{Q_{t}}$. If $\mathcal{O}$ is not chordless then replace it with a chordless cycle containing $v$ by Lemma 4.6. Because the subquiver of green vertices in $Q_{t}$ is acyclic, any chord between the green vertices $w_{i},w_{j}$ has the same orientation as the directed path it replaced. Therefore $\mathcal{O}$ is a chordless cycle which is not oriented. But the only edge on $\mathcal{O}$ which has a positive entry in $A_{t}$ is $w_{0}-v$, contradicting Theorem 2.26. So there are no cycles in $L_{v}(Q_{t})$ with a single arrow labeled $v$.
Suppose there is an oriented cycle in $L_{v}(Q_{t})$ with at least $2$ arrows labeled $v$, say
$w_{0}\stackrel{{\scriptstyle v}}{{\rightarrow}}w_{1}\rightarrow\cdots\rightarrow w_{j}\stackrel{{\scriptstyle v}}{{\rightarrow}}w_{j+1}\rightarrow\cdots\rightarrow w_{k}=w_{0}$
with $j>1$ and none of the arrows on the path $w_{1}\rightarrow w_{2}\rightarrow\cdots\rightarrow w_{j}$ labeled $v$ (the other arrows may or may not be labeled $v$). Thus $\mathcal{O}=(v-w_{1}-\cdots-w_{j}-v)$ is a cycle in $K_{Q_{t}}$ which is not oriented. After applying Lemma 4.6 if necessary, we may assume $\mathcal{O}$ is a chordless cycle. But the only edge on $\mathcal{O}$ which has a positive entry in $A_{t}$ is $w_{j}-v$, contradicting Theorem 2.26. Thus $L_{v}(Q_{t})$ is acyclic. ∎
Definition 5.6.
A linear extension of $L_{v}(Q_{t})$ is a linear order $<$ on the vertices such that $v_{i}<v_{j}$ whenever $v_{i}\rightarrow v_{j}$ (it does not matter whether the arrow is labeled $v$). Let $\sigma_{v}$ be a cyclic ordering associated to a linear extension of $L_{v}(Q_{t})$.
Example 5.7.
Continuing with Examples 5.2 and 5.4, there are several possible choices for the cyclic orderings $\sigma_{v}$ for each fixed $v$, though all of them result in wiggle equivalent COQs $(Q_{t},\sigma_{v})$. We could take
$\sigma_{v_{1}}=\sigma_{v_{2}}=\sigma_{v_{3}}=(v_{6},v_{1},v_{4},v_{3},v_{2},v_{5}),\text{ and}$
$\sigma_{v_{4}}=\sigma_{v_{5}}=\sigma_{v_{6}}=(v_{6},v_{3},v_{1},v_{4},v_{2},v_{5}).$
Lemma 5.8.
The matrix $U_{t}=(u_{v_{i}v_{j};t})$ is upper triangular when written with a linear extension $<$ of $L_{v}(Q_{t})$. Thus $U_{t}$ is a unipotent companion of $(Q_{t},\sigma_{v})$.
Proof.
Suppose $v_{i},v_{j}$ are vertices such that $u_{v_{i}v_{j};t}\neq 0$. We want to show that $v_{i}<v_{j}$. Because $|a_{v_{i}v_{j};t}|=|b_{v_{i}v_{j};t}|$, and $u_{v_{i}v_{j};t}=\frac{a_{v_{i}v_{j};t}-b_{v_{i}v_{j};t}}{2}\neq 0$, we have $a_{v_{i}v_{j};t}=-b_{v_{i}v_{j};t}$.
If $a_{v_{i}v_{j};t}>0$ then $v_{i}\leftarrow v_{j}$ (as $b_{v_{i}v_{j};t}<0$). By Definition 5.3, the graph $L_{v}(Q_{t})$ has an arrow ${v_{i}\rightarrow v_{j}}$. Thus $v_{i}<v_{j}$ in this case.
Otherwise $a_{v_{i}v_{j};t}<0$ and $v_{i}\rightarrow v_{j}$ in $L_{v}(Q_{t})$, so we have $v_{i}<v_{j}$. ∎
Example 5.9.
Continuing with Example 5.2, we find that $U_{t}$ is a unipotent companion of $(Q_{t},\sigma_{v})$, for any vertex $v$. As suggested by Example 5.7, if we use the linear order $v_{6}<v_{3}<v_{1}<v_{4}<v_{2}<v_{5}$, then
$U_{t}=\begin{pmatrix}1&-1&-1&0&1&0\\ 0&1&0&0&-1&1\\ 0&0&1&-1&0&0\\ 0&0&0&1&-1&0\\ 0&0&0&0&1&-1\\ 0&0&0&0&0&1\end{pmatrix}.$
Lemma 5.10.
The COQs $(Q_{t},\sigma_{v}),(Q_{t},\sigma_{w})$ are wiggle equivalent for all vertices $v,w$.
Proof.
By Lemma 5.8, the matrix $U_{t}$ is a unipotent companion of both $(Q_{t},\sigma_{v})$ and $(Q_{t},\sigma_{w})$. The claim follows from Lemma 4.7. ∎
Lemma 5.11.
The COQ $(Q_{t},\sigma_{v})$ is proper.
Proof.
It suffices to check that each vertex $q$ is proper in the COQ $(Q_{t},\sigma_{q})$ by Lemma 5.10. Let $<$ be a linear extension of $L_{q}(Q_{t})$. We consider all oriented paths ${p\rightarrow q\rightarrow r}$ in cases, based on the color (green or red) of $p,q,$ and $r$ in $Q_{t}$. As noted in Lemma 5.5, the subquiver of all red or all green vertices is acyclic. We have chosen our order $<$ to be compatible with these acyclic subquivers. Thus if all three of $p,q,r$ are green (resp., red) then $p\rightarrow q\rightarrow r$ is a right turn at $q$. A similar argument applies if $p$ or $r$ is a different color than the other two vertices.
This leaves us with the alternating patterns: red-green-red or green-red-green. By construction of $<$, if $p,r$ are both a different color than $q$, then there is an arrow $r\stackrel{{\scriptstyle q}}{{\rightarrow}}p$ in $L_{q}(Q_{t})$, and so $r<p$. Thus the path $p\rightarrow q\rightarrow r$ is again a right turn at $q$. ∎
Example 5.12.
Continuing with Examples 5.2 and 5.7, Figure 12 shows the COQ $(Q_{t},\sigma_{v_{1}})$. Note that every oriented path of length $2$ is a right turn.
A COQ whose quiver is as in the previous examples, and whose cyclic ordering is (v1, v4, v3, v2, v5, v6). Every vertex is proper, for example the oriented 2-path v2 to v6 to v1 is a right turn as v2 v6 v1 appear in clockwise order.
Figure 12. The proper COQ $(Q_{t},\sigma_{v_{1}})$.
Proof of Theorem 1.1.
Let $\sigma_{t}=\sigma_{v_{1}}$ be a cyclic ordering constructed from $Q_{t}$ as in Definition 5.6. By Lemma 5.8, $U_{t}$ is a unipotent companion of the COQ $(Q_{t},\sigma_{t})$. By Lemma 5.11, $(Q_{t},\sigma_{t})$ is proper. It remains to show that these COQs (as $t\in\mathbb{T}_{n}$ varies) are all proper mutation equivalent.
Fix an arbitrary $t^{\prime}\stackrel{{\scriptstyle i}}{{\text{---}}}t$. It suffices to show that the COQ $(\mu_{v_{i}}(Q_{t}),\sigma_{t^{\prime}})$ is in the wiggle equivalence class $\mu_{v_{i}}((Q_{t},\sigma_{t}))$. Let $<$ be a linear extension of $L_{v_{i}}(Q_{t})$. Say that $v_{i}$ is green (resp., red) in $Q_{t}$. For every vertex $w\in{\operatorname{Out}}(v_{i})$ (resp., ${\operatorname{In}}(v_{i})$), we have $v_{i}<w$ (resp. $w<v_{i}$). By Proposition 2.30, we have
$B_{t^{\prime}}=M_{Q_{t}}(v_{i},\varepsilon)B_{t}M_{Q_{t}}(v_{i},\varepsilon)^{T},$
$A_{t^{\prime}}=M_{Q_{t}}(v_{i},\varepsilon)A_{t}M_{Q_{t}}(v_{i},\varepsilon)^{T}$
for $\varepsilon=1$ (resp., $-1$). Thus
$U_{t^{\prime}}$ $=\frac{1}{2}\big{(}M_{Q_{t}}(v_{i},\varepsilon)A_{t}M_{Q_{t}}(v_{i},\varepsilon)^{T}-M_{Q_{t}}(v_{i},\varepsilon)B_{t}M_{Q_{t}}(v_{i},\varepsilon)^{T}\big{)}$
$=M_{Q_{t}}(v_{i},\varepsilon)U_{t}M_{Q_{t}}(v_{i},\varepsilon)^{T}.$
In particular, $U_{t^{\prime}}$ is exactly the unipotent companion of $\mu_{v_{i}}((Q_{t},\sigma_{t}))$ produced by Lemma 4.2 (resp. 4.1). The theorem then follows from Lemma 4.7. ∎
6. Applications
As noted in Theorem 3.16, the integral congruence class of a unipotent companion is a proper mutation invariant of COQs. If a COQ is totally proper, then every mutation is proper and so we have a mutation invariant of the underlying quiver.
Theorem 6.1 ([11, Theorem 11.3, Remark 11.11]).
A quiver $Q$ has at most one totally proper cyclic ordering (up to wiggle equivalence). There is an efficient algorithm that computes a cyclic ordering $\sigma_{Q}$ such that if $Q$ has a totally proper cyclic ordering, then $(Q,\sigma_{Q})$ is totally proper.
If a quiver $Q$ is mutation equivalent to a totally proper quiver $Q^{\prime}$, then the integral congruence class of a unipotent companion of the COQ $(Q,\sigma_{Q})$ will agree with that of $(Q^{\prime},\sigma_{Q^{\prime}})$. In particular, their Alexander polynomials and Markov invariants (Definition 3.19) will agree.
By Theorem 1.1, we now have many new totally proper quivers. Let us discuss a few examples.
Corollary 6.2.
Suppose $Q$ is an acyclic quiver whose underlying unoriented simple graph is a chordless cycle $\mathcal{O}$. Then a cyclic ordering $\sigma$ of the vertices of $Q$ is totally proper if and only if ${\operatorname{wind}}_{\sigma}(\mathcal{O})=0$.
Note that Corollary 6.2 makes no assumption about the multiplicities of the arrows.
Example 6.3.
As an illustration, we compute some mutation invariants for a few families. Let $\tilde{A}(r,\ell)$ denote a quiver with $n=r+\ell$ vertices $v_{1},\ldots,v_{n}$ which has $r$ locations with a single arrow $v_{i}\rightarrow v_{i+1}$ and $\ell$ locations with a single arrow $v_{i}\leftarrow v_{i+1}$ (see [13, Example 6.9]). While there are many ways to place the arrows, all are related by sink and/or source mutations. Thus we may assume that $\tilde{A}(r,\ell)$ has arrows
$v_{1}\rightarrow v_{2}\rightarrow\cdots\rightarrow v_{r}\leftarrow v_{r+1}\leftarrow\cdots\leftarrow v_{n}\leftarrow v_{1}.$
For $r,\ell>0$, the quiver $\tilde{A}(r,\ell)$ is acyclic, and thus totally proper. One totally proper cyclic ordering is $\sigma_{r\ell}=(v_{1},\ldots,v_{r-1},v_{n},v_{n-1},\ldots,v_{r+1},v_{r})$.
The Alexander polynomial of $(\tilde{A}(r,\ell),\sigma_{r\ell})$ is
$\Delta_{\tilde{A}(r,\ell)}(t)=(t^{\ell}-(-1)^{\ell})(t^{r}-(-1)^{r}).$
Every quiver mutation equivalent to $(\tilde{A}(r,\ell),\sigma_{r\ell})$ has integrally congruent unipotent companions, and in particular a matching Alexander polynomial.
Thus, if $r=3$ and $\ell=2$ then one unipotent companion is
$U=\begin{pmatrix}1&-1&0&-1&0\\ 0&1&-1&0&0\\ 0&0&1&0&-1\\ 0&0&0&1&-1\\ 0&0&0&0&1\end{pmatrix},$
and we can compute the Alexander polynomial
$\Delta_{\tilde{A}(3,2)}(t)=\det(tU-U^{T})=t^{5}-t^{3}+t^{2}-1.$
Example 6.4.
We can expand Example 6.3 by considering quivers whose simple undirected graphs are cycles but have multiple arrows between vertices. Choose nonnegative integers $r,\ell$ and positive integers $a_{1},\ldots,a_{r+\ell}$. Let $C(r,\ell)$ denote the quiver with $n=r+\ell$ vertices $v_{i}$, and arrows
$v_{1}\stackrel{{\scriptstyle a_{1}}}{{\rightarrow}}v_{2}\stackrel{{\scriptstyle a_{2}}}{{\rightarrow}}\cdots\stackrel{{\scriptstyle a_{r-1}}}{{\rightarrow}}v_{r}\stackrel{{\scriptstyle a_{r}}}{{\leftarrow}}v_{r+1}\stackrel{{\scriptstyle a_{r+1}}}{{\leftarrow}}\cdots\stackrel{{\scriptstyle a_{n-1}}}{{\leftarrow}}v_{n}\stackrel{{\scriptstyle a_{n}}}{{\leftarrow}}v_{1}.$
For $r,\ell>0$, the cyclic ordering $\sigma_{r\ell}$ is totally proper.
We compute the Alexander polynomials for several small values of $r,\ell$. Recall that for a $4$-vertex quiver $Q$ we can write the Alexander polynomial as
$\Delta_{Q}(t)=(t-1)^{4}+M_{Q}\cdot t(t-1)^{2}+\det(B_{Q})\cdot t^{2}.$
In particular,
$\Delta_{C(4,0)}(t)=$ $(t-1)^{4}+(a_{1}^{2}+a_{2}^{2}+a_{3}^{2}+a_{4}^{2}-a_{1}a_{2}a_{3}a_{4})t(t-1)^{2}$
$+(a_{1}^{2}a_{3}^{2}+a_{2}^{2}a_{4}^{2}-2a_{1}a_{2}a_{3}a_{4})t^{2};$
$\Delta_{C(3,1)}(t)=$ $(t-1)^{4}+(a_{1}^{2}+a_{2}^{2}+a_{3}^{2}+a_{4}^{2}+a_{1}a_{2}a_{3}a_{4})t(t-1)^{2}$
$+(a_{1}^{2}a_{3}^{2}+a_{2}^{2}a_{4}^{2}+2a_{1}a_{2}a_{3}a_{4})t^{2};$
$\Delta_{C(2,2)}(t)=$ $(t-1)^{4}+(a_{1}^{2}+a_{2}^{2}+a_{3}^{2}+a_{4}^{2})t(t-1)^{2}$
$+(a_{1}^{2}a_{3}^{2}+a_{2}^{2}a_{4}^{2}-2a_{1}a_{2}a_{3}a_{4})t^{2}.$
For a $5$-vertex quiver $Q$, we can write the Alexander polynomial as
$\Delta_{Q}(t)=(t-1)^{5}+M_{Q}\cdot t(t-1)^{3}+d\cdot t^{2}(t-1)$
for some integer coefficient $d$. In particular,
$\Delta_{C(4,1)}(t)=$ $(t-1)^{5}+(a_{1}^{2}+a_{2}^{2}+a_{3}^{2}+a_{4}^{2}+a_{5}^{2}+a_{1}a_{2}a_{3}a_{4}a_{5})t(t-1)^{3}$
$+(a_{1}^{2}a_{3}^{2}+a_{1}^{2}a_{4}^{2}+a_{2}^{2}a_{4}^{2}+a_{2}^{2}a_{5}^{2}+a_{3}^{2}a_{5}^{2}-3a_{1}a_{2}a_{3}a_{4}a_{5})t^{2}(t-1);$
$\Delta_{C(3,2)}(t)=$ $(t-1)^{5}+(a_{1}^{2}+a_{2}^{2}+a_{3}^{2}+a_{4}^{2}+a_{5}^{2})t(t-1)^{3}$
$+(a_{1}^{2}a_{3}^{2}+a_{1}^{2}a_{4}^{2}+a_{2}^{2}a_{4}^{2}+a_{2}^{2}a_{5}^{2}+a_{3}^{2}a_{5}^{2}-a_{1}a_{2}a_{3}a_{4}a_{5})t^{2}(t-1).$
These examples suggest a pattern for the Markov invariant, agreeing with Proposition 3.21. Without loss of generality, assume $r>0$. Then
$M_{C(r,\ell)}=\begin{cases}\sum_{i}^{n}a_{i}^{2}&\text{if }\ell\geq 2;\\ \sum_{i}^{n}a_{i}^{2}+\prod_{i}^{n}a_{i}&\text{if }\ell=1;\\ \sum_{i}^{n}a_{i}^{2}-\prod_{i}^{n}a_{i}&\text{if }\ell=0.\end{cases}$
Remark 6.5.
Recall Lagrange’s four squares theorem, which states that every nonnegative integer is the sum of at most four squares (of integers). Thus there is a totally proper $4$-vertex COQ with Markov invariant $x$ for every $x\geq 0$. However some integers are not the sum of exactly four positive squares. For example, $2^{2k+1}$, or $29$ (see [23]). Thus any quiver whose associated Markov invariant is one of these values must not be mutation equivalent to $C(2,2)$ (for any positive values $a_{i}$).
We have the following corollaries of Proposition 3.21.
Corollary 6.6.
Let $Q$ be a connected, totally proper and mutation-acyclic COQ on $n$ vertices. Then the Markov invariant $M_{Q}\geq n-1$. Further, if $M_{Q}=n-1$ then $Q$ is mutation equivalent to a quiver whose underlying undirected graph is a tree (with no parallel arrows).
Corollary 6.7.
Let $Q$ be a totally proper and mutation-acyclic COQ on $n$ vertices. Then the Markov invariant and the determinant of the exchange matrix satisfy the following inequality:
$\det(B_{Q})\leq(2M_{Q})^{n/2}.$
Proof.
It suffices to consider acyclic COQs $Q$ with vertices $v_{1},\ldots,v_{n}$. By Proposition 3.21, $\sum_{i<j}b_{v_{i}v_{j}}^{2}\leq M_{Q}.$ Recall the Frobenius norm $||B_{Q}||_{F}=\big{(}\sum_{i,j}b_{v_{i}v_{j}}^{2}\big{)}^{1/2}$ of $B_{Q}$ (see also the Hilbert-Schmidt norm or Schur norm), thus $||B_{Q}||_{F}^{2}\leq 2M_{Q}.$
As $\det(B_{Q})=0$ if $n$ is odd, we assume $n$ is even. As $B_{Q}$ is skew-symmetric, its eigenvalues are all imaginary and come in signed pairs, say they have norms $|\lambda_{1}|,\ldots,|\lambda_{n/2}|$. Let $\lambda=\max(|\lambda_{j}|)$. Then $\det(B_{Q})\leq\lambda^{n}$. As $B_{Q}$ is skew-symmetric the spectral norm $\rho(B_{Q})=\lambda$. It is a classical fact that the spectral norm is a lower bound on the Frobenius norm: $\lambda\leq||B_{Q}||_{F}$. Therefore we have
$\det(B_{Q})\leq\lambda^{n}\leq||B_{Q}||_{F}^{n}\leq(2M_{Q})^{n/2}.\qed$
Corollaries 6.6 and 6.7 give short proofs that certain quivers are not mutation-acyclic.
Example 6.8.
The Somos sequences are integer sequences, the first $5$ of which are associated to cluster algebras with particularly symmetric quivers. The Somos-$4$ quiver $S$ is shown in Figure 13. The distinguished cyclic ordering $\sigma_{S}=(v_{1},v_{4},v_{2},v_{3})$ (see [11][Remark 11.11]) is the only potentially totally proper ordering for $S$. This quiver is not mutation-acyclic though, which is quickly determined by computing the Markov invariant of the COQ $(S,\sigma_{S})$ and applying Corollary 6.6:
$M_{S}=1^{2}+1^{2}+2^{2}+3^{2}+2^{2}+1^{2}-6-6+2+2-12=0.$
A four vertex quiver with vertices v1, v2, v3, v4 and arrows as follows: One arrow from v1 to v2, v3 to v4, and v1 to v4; two arrows from v3 to v1 and v4 to v2; three arrows from v2 to v3. Mutation at v1 gives an isomorphic quiver.
Figure 13. The Somos-$4$ quiver $S$.
Remark 6.9.
Fix an integer $a\geq 1$. Example 6.8 generalizes to the $4$-vertex quiver $aS$, which has the same vertices as $S$, and $a$ arrows for every arrow in $S$ (thus, for example, there are $3a$ arrows from $v_{2}$ to $v_{3}$). Each has a nonpositive Markov invariant.
Example 6.10.
Continuing with Example 6.4, consider the COQ $C(4,0)$ with cyclic ordering $(v_{1},v_{2},v_{3},v_{4})$, shown below.
A COQ with cyclic ordering (v1, v2, v3, v4) and a1 arrows from v1 to v2, a2 arrows from v2 to v3, a3 arrows from v3 to v4, and a4 arrows from v4 to v1.
Some such quivers are mutation-acyclic, for example when $a_{i}=1$ for all $i$ then $C(4,0)$ has type $D_{4}$. Others are not. Warkentin showed in [26][Example 6.11] that if $a_{1}=a_{3}\geq 2$ and $a_{2}=a_{4}\geq 2$ then this quiver is not mutation-acyclic (by a precise description of the mutation class). The same can be observed by Corollary 6.6, noting that $M_{C(4,0)}\leq 0$ for such quivers.
Suppose that $a_{4}=a_{1}a_{2}a_{3}$. Then $M_{C(4,0)}=a_{1}^{2}+a_{2}^{2}+a_{3}^{2}\geq 3.$ Thus we cannot determine whether the quiver is mutation-acylic or not from Corollary 6.6. However,
$\det(B_{C(4,0)})=(a_{1}a_{3}+a_{1}a_{2}^{2}a_{3})^{2}=(a_{2}^{2}+1)^{2}(a_{1}a_{3})^{2}.$
So we have $\det(B_{C(4,0)})>(2M_{C(4,0)})^{2}$ whenever (for example) $a_{2}\geq\max(a_{1},a_{3})$ and $a_{1}a_{3}\geq 6$. Thus, for these values of $a_{i}$, the quiver $C(4,0)$ is again not mutation-acyclic by Corollary 6.7.
Example 6.11.
Consider the quivers associated with the elliptic root system of double extended type $E_{7}$ or $E_{8}$, denoted $E_{7}^{(1,1)}$ or $E_{8}^{(1,1)}$ and shown in Figure 14. (Note that the quiver in Figure 1 is mutation equivalent to $E_{8}^{(1,1)}$.) These quivers have finite mutation classes, and can be shown to be totally proper by exhaustive computation. (We emphasize the exhaustion; $E_{8}^{(1,1)}$ has $5739$ quivers, up to isomorphism, in its class.) These quivers are not mutation-acyclic. This has been shown by brute force, or by representation theory [17][Section 2.3]. We sketch a new proof.
The canonical candidate cyclic orderings $\sigma_{E_{7}^{(1,1)}}$ and $\sigma_{E_{8}^{(1,1)}}$ for the quivers shown in Figure 1 are $(v_{1},\ldots,v_{9})$ and $(v_{1},\ldots,v_{10})$ respectively. (While $(E_{7}^{(1,1)},\sigma_{E_{7}^{(1,1)}})$ and $(E_{8}^{(1,1)},\sigma_{E_{8}^{(1,1)}})$ are totally proper, we do not rely on this fact.) So we treat $E_{7}^{(1,1)}$ and $E_{8}^{(1,1)}$ as COQs with their respective cyclic ordering. Their Markov invariants are:
$M_{E_{7}^{(1,1)}}=10+4-6=8,\quad M_{E_{8}^{(1,1)}}=11+4-6=9.$
By Corollary 6.6, the only acyclic quivers that $E_{7}^{(1,1)}$ (resp., $E_{8}^{(1,1)}$) could be mutation equivalent to are tree quivers (with no parallel arrows).
Up to isomoprhism, there are $47$ (unoriented) trees with $9$ vertices and $106$ trees with $10$ vertices. All orientations of each of these trees are mutation equivalent. By Theorem 3.11, all cyclic orderings of a tree are wiggle equivalent. Thus any quiver mutation equivalent to a tree on $9$ (resp., $10$) vertices must have one of the $47$ (resp., $106$) Alexander polynomials. (Actually the list is slightly smaller, as some non-isomorphic trees have the same Alexander polynomial.)
So it suffices to check if $\Delta_{E_{7}^{(1,1)}}(t)$ and $\Delta_{E_{8}^{(1,1)}}(t)$ appear in these lists. A (still tedious, but straightforward and human verifiable) computation shows that they are not. The computation can be further simplified using the work of Amanda Schwartz, see Remark 3.23.
A nine vertex quiver with vertices v1, ..., v9 and arrows as follows: 2 arrows from v5 to v1, one arrow from v1 to v2, v2 to v5, v1 to v3, v3 to v5, v1 to v4, v4 to v5, v3 to v6, v6 to v7, v4 to v8, and v8 to v9. Thus there are oriented three cycles on v5, v1, and any one of v2, v3, or v4.
A ten vertex quiver with vertices v1, ..., v10 and arrows as follows: 2 arrows from v5 to v1, one arrow from v1 to v2, v2 to v5, v1 to v3, v3 to v5, v1 to v4, v4 to v5, v3 to v6, v6 to v7, v7 to v8, v8 to v9, and v4 to v10. Thus there are oriented three cycles on v5, v1, and any one of v2, v3, or v4.
Figure 14. The extended affine quivers $E_{7}^{(1,1)}$ and $E_{8}^{(1,1)}$.
Example 6.12.
Choose positive integers $a,b,c$. Let $Q$ be the $4$-vertex COQ below:
A four vertex COQ with cyclic ordering (v1, v2, v3, v4) and arrows as follows: a arrows from v1 to v2; b arrows from v2 to v3 and from v2 to v4; c arrows from v3 to v1 and v4 to v1.
Then $M_{Q}=a^{2}+2b^{2}+2c^{2}-2abc.$ Let $Q^{\prime}$ be the subquiver supported by $v_{1},v_{2},v_{3}$, thus $M_{Q^{\prime}}=a^{2}+b^{2}+c^{2}-abc.$ Note that $M_{Q}=2M_{Q^{\prime}}-a^{2}.$ Thus $M_{Q}<0$ whenever $M_{Q^{\prime}}<\frac{a^{2}}{2}$. (This last inequality is generally satisfied for mutation-infinite quivers; as $M_{Q^{\prime}}$ is constant on the mutation class, but the number of arrows grows arbitrarily large, we can find many quivers with some weight $a$ much larger than $M_{Q^{\prime}}$.) Thus by Corollary 6.6, such $Q$ cannot be mutation-acyclic.
Example 6.13.
We next illustrate a way of combining Corollary 6.6 with another mutation invariant [20], the multiset of GCDs of all the integers in the columns (or rows) of the $B$-matrix. We construct a COQ as in Example 6.12 with $a=b=2$ and $c=4$. Thus $M_{Q}=12\geq 3$, and we cannot use Corollary 6.6 to conclude that $Q$ is not mutation-acyclic. However, note that the GCD of the entries of $B_{Q}$ is $2$. Thus if $Q$ is mutation-acyclic, it must be mutation equivalent to a connected acyclic quiver with the same GCD and $M_{Q}=12$. There are two such quivers, both trees with $2$ arrows between each pair of neighboring vertecies. The $B$-matrices of these acyclic quivers have determinant 00 and $16$, but $\det(B_{Q})=144$. Thus $Q$ is not mutation-acyclic.
Finally, we note that Lemmas 4.1 4.2 extend to the admissible quasi-Cartan companion $A_{U}=U+U^{T}$ (discussed in [11][Remark 13.2]).
Corollary 6.14.
Fix a totally proper COQ $Q$ and a proper vertex $v_{j}$. Let $U_{Q}$ be a unipotent companion with linear order $<$ so that $v_{i}<v_{j}$ (resp., $v_{j}<v_{i}$) for all $v_{i}\in{\operatorname{In}}(v_{j})$ (resp., ${\operatorname{Out}}(v_{j})$). Then $M_{Q}(v_{j},\epsilon)(U_{Q}+U_{Q}^{T})M_{Q}(v_{j},\epsilon)$ is an admissible quasi-Cartan companion of $\mu_{v_{j}}(Q)$ when $\epsilon=-1$ (resp., $1$).